### v7

# A Comprehensive Overview of Large Language Models

Abstract
Large Language Models (LLMs) have recently demonstrated remarkable capabilities in natural language processing tasks and beyond. This success of LLMs has led to a large influx of research contributions in this direction. These works encompass diverse topics such as architectural innovations, better training strategies, context length improvements, fine-tuning, multi-modal LLMs, robotics, datasets, benchmarking, efficiency, and more. With the rapid development of techniques and regular breakthroughs in LLM research, it has become considerably challenging to perceive the bigger picture of the advances in this direction. Considering the rapidly emerging plethora of literature on LLMs, it is imperative that the research community is able to benefit from a concise yet comprehensive overview of the recent developments in this field. This article provides an overview of the existing literature on a broad range of LLM-related concepts. Our self-contained comprehensive overview of LLMs discusses relevant background concepts along with covering the advanced topics at the frontier of research in LLMs. This review article is intended to not only provide a systematic survey but also a quick comprehensive reference for the researchers and practitioners to draw insights from extensive informative summaries of the existing works to advance the LLM research.
Index Terms:
Large Language Models, LLMs, chatGPT, Augmented LLMs, Multimodal LLMs, LLM training, LLM Benchmarking
Figure 1: The trends in the number of LLM models introduced over the years.
Figure 2: Chronological display of LLM releases: light blue rectangles represent ‘pre-trained’ models, while dark rectangles correspond to ‘instruction-tuned’ models. Models on the upper half signify open-source availability, whereas those on the bottom half are closed-source. The chart illustrates the increasing trend towards instruction-tuned models and open-source models, highlighting the evolving landscape and trends in natural language processing research.
I Introduction
Language plays a fundamental role in facilitating communication and self-expression for humans, and their interaction with machines. The need for generalized models stems from the growing demand for machines to handle complex language tasks, including translation, summarization, information retrieval, conversational interactions, etc. Recently, significant breakthroughs have been witnessed in language models, primarily attributed to transformers [1], increased computational capabilities, and the availability of large-scale training data. These developments have brought about a revolutionary transformation by enabling the creation of LLMs that can approximate human-level performance on various tasks [2, 3]. Large Language Models (LLMs) have emerged as cutting-edge artificial intelligence systems that can process and generate text with coherent communication [4], and generalize to multiple tasks [5, 6].
The historical progress in natural language processing (NLP) evolved from statistical to neural language modeling and then from pre-trained language models (PLMs) to LLMs. While conventional language modeling (LM) trains task-specific models in supervised settings, PLMs are trained in a self-supervised setting on a large corpus of text [7, 8, 9] with the aim to learn generic representation shareable among various NLP tasks. After fine-tuning for downstream tasks, PLMs surpass the performance gains of traditional language modeling (LM). The larger PLMs bring more performance gains, which has led to the transitioning of PLMs to LLMs by significantly increasing model parameters (tens to hundreds of billions) [10] and training dataset (many GBs and TBs) [10, 11]. Following this development, numerous LLMs have been proposed in the literature [10, 11, 12, 6, 13, 14, 15]. An increasing trend in the number of released LLMs and names of a few significant LLMs proposed over the years are shown in Fig 1 and Fig 2, respectively.
The early work on LLMs, such as T5 [10] and mT5 [11] employed transfer learning until GPT-3 [6] showing LLMs are zero-shot transferable to downstream tasks without fine-tuning. LLMs accurately respond to task queries when prompted with task descriptions and examples. However, pre-trained LLMs fail to follow user intent and perform worse in zero-shot settings than in few-shot. Fine-tuning them with task instructions data [16, 17, 18, 19] and aligning with human preferences [20, 21] enhances generalization to unseen tasks, improving zero-shot performance significantly and reducing misaligned behavior.
Additional to better generalization and domain adaptation, LLMs appear to have emergent abilities, such as reasoning, planning, decision-making, in-context learning, answering in zero-shot settings, etc. These abilities are known to be acquired by them due to their gigantic scale even when the pre-trained LLMs are not trained specifically to possess these attributes [22, 23, 24]. Such abilities have led LLMs widely adopted in diverse settings including, multi-modal, robotics, tool manipulation, question answering, autonomous agents, etc. Various improvements have also been suggested in these areas either by task-specific training [25, 26, 27, 28, 29, 30, 31] or better prompting [32].
Figure 3: A broader overview of LLMs, dividing LLMs into five branches: 1. Training 2. Inference 3. Evaluation 4. Applications 5. Challenges
The LLMs abilities to solve diverse tasks with human-level performance come at a cost of slow training and inference, extensive hardware requirements, and higher running costs. Such requirements have limited their adoption and opened up opportunities to devise better architectures [15, 33, 34, 35] and training strategies [36, 37, 21, 38, 39, 40, 41]. Parameter efficient tuning [38, 41, 40], pruning, quantization, knowledge distillation, and context length interpolation [42, 43, 44, 45] among others are some of the methods widely studied for efficient LLM utilization.
Due to the success of LLMs on a wide variety of tasks, the research literature has recently experienced a large influx of LLM-related contributions. Researchers have organized the LLMs literature in surveys [46, 47, 48, 49], and topic-specific surveys in [50, 51, 52, 53, 54]. In contrast to these surveys, our contribution focuses on providing a comprehensive yet concise overview of the general direction of LLM research. This article summarizes architectural and training details of pre-trained LLMs and delves deeper into the details of concepts like fine-tuning, multi-modal LLMs, robotics, augmented LLMs, datasets, evaluation, and others to provide a self-contained comprehensive overview. Our key contributions are summarized as follows.
  • We present a survey on the developments in LLM research with the specific aim of providing a concise yet comprehensive overview of the direction.
  • We present extensive summaries of pre-trained models that include fine-grained details of architecture and training details.
  • Besides paying special attention to the chronological order of LLMs throughout the article, we also summarize major findings of the popular contributions and provide detailed discussion on the key design and development aspects of LLMs to help practitioners to effectively leverage this technology.
  • In this self-contained article, we cover a range of concepts to comprehend the general direction of LLMs comprehensively, including background, pre-training, fine-tuning, robotics, multi-modal LLMs, augmented LLMs, datasets, evaluation, etc.
We loosely follow the existing terminologies to ensure providing a more standardized outlook of this research direction. For instance, following [46], our survey discusses pre-trained LLMs with 10B parameters or more. We refer the readers interested in smaller pre-trained models to [47, 48, 49].
The organization of this paper is as follows. Section II discusses the background of LLMs. Section III focuses on LLMs overview, architectures, training pipelines and strategies, and utilization in different aspects. Section IV presents the key findings derived from each LLM. Section V highlights the configuration and parameters that play a crucial role in the functioning of these models. Summary and discussions are presented in section VIII. The LLM training and evaluation, datasets and benchmarks are discussed in section VI, followed by challenges and future directions and conclusion in sections IX and X, respectively.
II Background
We provide the relevant background to understand the fundamentals related to LLMs in this section. Aligned with our objective of providing a comprehensive overview of this direction, this section offers a comprehensive yet concise outline of the basic concepts. We focus more on the intuitive aspects and refer the readers interested in details to the original works.
II-A Tokenization
LLMs are trained on text to predict text, and similar to other natural language processing systems, they use tokenization [55] as the essential preprocessing step. It aims to parse the text into non-decomposing units called tokens. Tokens can be characters, subwords [56], symbols [57], or words, depending on the size and type of the model. Some of the commonly used tokenization schemes in LLMs are briefed here. Readers are encouraged to refer to [58] for a detailed survey.
II-A1 WordPiece [59]
It was introduced in [59] as a novel text segmentation technique for Japanese and Korean languages to improve the language model for voice search systems. WordPiece selects tokens that increase the likelihood of an n-gram-based language model trained on the vocabulary composed of tokens.
II-A2 BPE [57]
Byte Pair Encoding (BPE) has its origin in compression algorithms. It is an iterative process of generating tokens where pairs of adjacent symbols are replaced by a new symbol, and the occurrences of the most occurring symbols in the input text are merged.
II-A3 UnigramLM [56]
In this tokenization, a simple unigram language model (LM) is trained using an initial vocabulary of subword units. The vocabulary is pruned iteratively by removing the lowest probability items from the list, which are the worst performing on the unigram LM.
II-B Attention
Attention, particularly selective attention, has been widely studied under perception, psychophysics, and psychology. Selective attention can be conceived as “the programming by the O of which stimuli will be processed or encoded and in what order this will occur” [60]. While this definition has its roots in visual perception, it has uncanny similarities with the recently formulated attention [61, 62] (which stimuli will be processed) and positional encoding (in what order this will occur) [62] in LLMs. We discuss both in sections II-C and II-D, respectively.
II-C Attention in LLMs
The attention mechanism computes a representation of the input sequences by relating different positions (tokens) of these sequences. There are various approaches to calculating and implementing attention, out of which some famous types are given below.
II-C1 Self-Attention [62]
The self-attention is also known as intra-attention since all the queries, keys, and values come from the same block (encoder or decoder). The self-attention layer connects all the sequence positions with $O(1)$ space complexity which is highly desirable for learning long-range dependencies in the input.
II-C2 Cross Attention
In encoder-decoder architectures, the outputs of the encoder blocks act as the queries to the intermediate representation of the decoder, which provides the keys and values to calculate a representation of the decoder conditioned on the encoder. This attention is called cross-attention.
II-C3 Full Attention
The naive implementation of calculating self-attention is known as full attention.
II-C4 Sparse Attention [63]
The self-attention has a time complexity of $O(n^{2})$, which becomes prohibitive when scaling the LLMs to large context windows. An approximation to the self-attention was proposed in [63], which greatly enhanced the capacity of GPT series LLMs to process a greater number of input tokens in a reasonable time.
II-C5 Flash Attention [64]
The bottleneck for calculating the attention using GPUs lies in the memory access rather than the computational speed. Flash Attention uses the classical input tiling approach to process the blocks of the input in GPU on-chip SRAM rather than doing IO for every token from the High Bandwith Memory (HBM). An extension of this approach to sparse attention follows the speed gains of the full attention implementation. This trick allows even greater context-length windows in the LLMs as compared to those LLMs with sparse attention.
II-D Encoding Positions
The attention modules do not consider the order of processing by design. Transformer [62] introduced “positional encodings” to feed information about the position of the tokens in input sequences. Several variants of positional encoding have been proposed [65, 66]. Interestingly, a recent study [67] suggests that adding this information may not matter for the state-of-the-art decoder-only Transformers.
II-D1 Absolute
This is the most straightforward approach to adding the sequence order information by assigning a unique identifier to each position of the sequence before passing it to the attention module.
II-D2 Relative
To pass the information on the relative dependencies of different tokens appearing at different locations in the sequence, a relative positional encoding is calculated by some kind of learning. Two famous types of relative encodings are:
Alibi: [65] In this approach, a scalar bias is subtracted from the attention score calculated using two tokens which increases with the distance between the positions of the tokens. This learned approach effectively favors using recent tokens for attention.
RoPE: Keys, queries, and values are all vectors in the LLMs. RoPE [66] involves the rotation of the query and key representations at an angle proportional to their absolute positions of the tokens in the input sequence. This step results in a relative positional encoding scheme which decays with the distance between the tokens.
II-E Activation Functions
The activation functions serve a crucial role in the curve-fitting abilities of the neural networks, as proved in [68]. The modern activation functions used in LLMs are different from the earlier squashing functions but are critical to the success of LLMs. We discuss these activation functions in this section.
II-E1 ReLU [69]
Rectified linear unit (ReLU) is defined as
$ReLU(x)=max(0,x)$ (1)
II-E2 GeLU [70]
Gaussian Error Linear Unit (GeLU) is the combination of ReLU, dropout [71] and zoneout [72]. It is the most widely used activation function in contemporary LLM literature.
II-E3 GLU variants [73]
Gated Linear Unit [74] is a neural network layer that is an element-wise product ($\otimes$) of a linear transformation and a sigmoid transformed ($\sigma$) linear projection of the input given as
$GLU(x,W,V,b,c)=(xW+b)\otimes\sigma(xV+c),$ (2)
where $X$ is the input of layer and $l$, $W,b,V\textnormal{ and }c$ are learned parameters.
GLU was modified in [73] to evaluate the effect of different variations in the training and testing of transformers, resulting in better empirical results. Here are the different GLU variations introduced in [73] and used in LLMs.
$ReGLU(x,W,V,b,c)$ $=max(0,xW+b)\otimes,$
$GEGLU(x,W,V,b,c)$ $=GELU(xW+b)\otimes(xV+c),$
$SwiGLU(x,W,V,b,c,\beta)$ $=Swish\beta(xW+b)\otimes(xV+c).$
II-F Layer Normalization
Layer normalization leads to faster convergence and is a widely used component in transformers. In this section, we provide different normalization techniques widely used in LLM literature.
II-F1 LayerNorm
Layer norm computes statistics over all the hidden units in a layer $(l)$ as follows:
$u^{l}=\frac{1}{n}\sum_{i}^{n}a_{i}^{l}\hskip 20.00003pt\sigma^{l}=\sqrt{\frac{1}{n}\sum_{i}^{n}(a_{i}^{l}-u^{l})^{2}},$ (3)
where $n$ is the number of neurons in the layer $l$ and $a_{i}^{l}$ is the summed input of the $i$ neuron in layer $l$. LayerNorm provides invariance to rescaling of the weights and re-centering of the distribution.
II-F2 RMSNorm
[75] proposed that the invariance properties of LayerNorm are spurious, and we can achieve the same performance benefits as we get from LayerNorm by using a computationally efficient normalization technique that trades off re-centering invariance with speed. LayerNorm gives the normalized summed input to layer $l$ as follows
$\overline{a_{i}^{l}}=\frac{a_{i}^{l}-u^{l}}{\sigma}g_{i}^{l}$ (4)
where $g_{i}^{l}$ is the gain parameter. RMSNorm [75] modifies $\overline{a_{i}^{l}}$ as
$\overline{a_{i}^{l}}=\frac{a_{i}^{l}}{\textnormal{RMS}(\mathbf{a}^{l})}g_{i}^{l},\hskip 3.00003pt\textnormal{where}\hskip 3.00003pt\textnormal{RMS}(\mathbf{a}^{l})=\sqrt{\frac{1}{n}\sum_{i}^{n}(a_{i}^{l})^{2}}.$ (5)
II-F3 Pre-Norm and Post-Norm
LLMs use transformer [62] architecture with some variations. The original implementation [62] used layer normalization after the residual connection, commonly called post-LN, concerning the order of Multihead attention – Residual – LN. There is another order of the normalization, referred to as pre-LN [76] due to the position of the normalization step before the self-attention layer as in LN – Multihead attention – Residual. Pre-LN is known to provide more stability in the training [77].
II-F4 DeepNorm
While pre-LN has certain benefits over post-LN training, pre-LN training has an unwanted effect on the gradients [77]. The earlier layers have larger gradients than those at the bottom. DeepNorm [78] mitigates these adverse effects on the gradients. It is given as
$\mathbf{x}^{l_{f}}=LN(\alpha\mathbf{x}^{l_{p}}+G^{l_{p}}(\mathbf{x}^{l_{p}},{\boldmath\theta}{{}^{l_{p}}}),$ (6)
where $\alpha$ is a constant and $\theta^{l_{p}}$ represents the parameters of layer $l_{p}$. These parameters are scaled by another constant $\beta$. Both of these constants depend only on the architecture.
II-G Distributed LLM Training
This section describes distributed LLM training approaches briefly. More details are available in [13, 37, 79, 80].
II-G1 Data Parallelism
Data parallelism replicates the model on multiple devices where data in a batch gets divided across devices. At the end of each training iteration weights are synchronized across all devices.
II-G2 Tensor Parallelism
Tensor parallelism shards a tensor computation across devices. It is also known as horizontal parallelism or intra-layer model parallelism.
II-G3 Pipeline Parallelism
Pipeline parallelism shards model layers across different devices. This is also known as vertical parallelism.
II-G4 Model Parallelism
A combination of tensor and pipeline parallelism is known as model parallelism.
II-G5 3D Parallelism
A combination of data, tensor, and model parallelism is known as 3D parallelism.
II-G6 Optimizer Parallelism
Optimizer parallelism also known as zero redundancy optimizer [37] implements optimizer state partitioning, gradient partitioning, and parameter partitioning across devices to reduce memory consumption while keeping the communication costs as low as possible.
II-H Libraries
Some commonly used libraries for LLMs training are: 1) Transformers [81], 2) DeepSpeed [36], 3) Megatron-LM [79], 4) JAX [82], 5) Colossal-AI [83], 6) BMTrain [80], 7) FastMoE [84], and frameworks are 1) MindSpore [85], 2) PyTorch [86], 3) Tensorflow [87], 4) MXNet [88].
II-I Data PreProcessing
This section briefly summarizes data preprocessing techniques used in LLMs training.
II-I1 Quality Filtering
For better results, training data quality is essential. Some approaches to filtering data are: 1) classifier-based and 2) heuristics-based. Classifier-based approaches train a classifier on high-quality data and predict the quality of text for filtering, whereas heuristics-based employ some rules for filtering like language, metrics, statistics, and keywords.
II-I2 Data Deduplication
Duplicated data can affect model performance and increase data memorization; therefore, to train LLMs, data deduplication is one of the preprocessing steps. This can be performed at multiple levels, like sentences, documents, and datasets.
II-I3 Privacy Reduction
Most of the training data for LLMs is collected through web sources. This data contains private information; therefore, many LLMs employ heuristics-based methods to filter information such as names, addresses, and phone numbers to avoid learning personal information.
II-J Architectures
Here we discuss the variants of the transformer architectures at a higher level which arise due to the difference in the application of the attention and the connection of transformer blocks. An illustration of attention patterns of these architectures is shown in Figure 4.
II-J1 Encoder Decoder
Transformers were originally designed as sequence transduction models and followed other prevalent model architectures for machine translation systems. They selected encoder-decoder architecture to train human language translation tasks. This architecture is adopted by [10, 89]. In this architectural scheme, an encoder encodes the input sequences to variable length context vectors, which are then passed to the decoder to maximize a joint objective of minimizing the gap between predicted token labels and the actual target token labels.
II-J2 Causal Decoder
The underlying objective of an LLM is to predict the next token based on the input sequence. While additional information from the encoder binds the prediction strongly to the context, it is found in practice that the LLMs can perform well in the absence of encoder [90], relying only on the decoder. Similar to the original encoder-decoder architecture’s decoder block, this decoder restricts the flow of information backward, i.e., the predicted token $t_{k}$ only depends on the tokens preceded by and up to $t_{k-1}$. This is the most widely used variant in the state-of-the-art LLMs.
II-J3 Prefix Decoder
The causal masked attention is reasonable in the encoder-decoder architectures where the encoder can attend to all the tokens in the sentence from every position using self-attention. This means that the encoder can also attend to tokens $t_{k+1}$ to $t_{n}$ in addition to the tokens from $t_{1}$ to $t_{k-1}$ while calculating the representation for $t_{k}$. But when we drop the encoder and only keep the decoder, we also lose this flexibility in attention. A variation in the decoder-only architectures is by changing the mask from strictly causal to fully visible on a portion of the input sequence, as shown in Figure 4. The Prefix decoder is also known as non-causal decoder architecture.
Figure 4: An example of attention patterns in language models, image is taken from [91].
Figure 5: An example of language model training objectives, image from [91].
II-K Pre-Training Objectives
This section describes LLMs pre-training objectives. For more details see the paper [91].
Figure 6: A basic flow diagram depicting various stages of LLMs from pre-training to prompting/utilization. Prompting LLMs to generate responses is possible at different training stages like pre-training, instruction-tuning, or alignment tuning.
II-K1 Full Language Modeling
An autoregressive language modeling objective where the model is asked to predict future tokens given the previous tokens, an example is shown in Figure 5.
II-K2 Prefix Language Modeling
A non-causal training objective, where a prefix is chosen randomly and only remaining target tokens are used to calculate the loss. An example is shown in Figure 5.
II-K3 Masked Language Modeling
In this training objective, tokens or spans (a sequence of tokens) are masked randomly and the model is asked to predict masked tokens given the past and future context. An example is shown in Figure 5.
II-K4 Unified Language Modeling
Unified language modeling [92] is a combination of causal, non-causal, and masked language training objectives. Here in masked language modeling, the attention is not bidirectional but unidirectional, attending either left-to-right or right-to-left context.
II-L Model Adaptation
This section discusses the fundamentals of LLMs adaptation stages, from pre-training to fine-tuning for downstream tasks and utilization. An example of different training stages and inference in LLMs is shown in Figure 6. In this paper, we refer alignment-tuning to aligning with human preferences, while occasionally the literature uses the term alignment for different purposes.
II-L1 Pre-Training
In the very first stage, the model is trained in a self-supervised manner on a large corpus to predict the next tokens given the input. The design choices of LLMs vary from encoder-decoder to decoder-only architectures with different building blocks and loss functions in sections II-FII-E,  II-K.
II-L2 Fine-Tuning
There are different styles to fine-tune an LLM. This section briefly discusses fine-tuning approaches.
  Transfer Learning: The pre-trained LLMs perform well for various tasks [6, 15]. But to improve the performance for a downstream task, pre-trained models are fine-tuned with the task-specific data [10, 11], known as transfer learning.   Instruction-tuning: To enable a model to respond to user queries effectively, the pre-trained model is fine-tuned on instruction formatted data i.e., instruction and an input-output pair. Instructions generally comprise multi-task data in plain natural language, guiding the model to respond according to the prompt and the input. This type of fine-tuning improves zero-shot generalization and downstream task performance. Details on formatting instruction data and its various styles are available in [16, 46, 93].
  Alignment-tuning: LLMs are prone to generate false, biased, and harmful text. To make them helpful, honest, and harmless models are aligned using human feedback. Alignment involves asking LLMs to generate unexpected responses and then updating their parameters to avoid such responses [20, 21, 94].
It ensures LLMs operate according to human intentions and values. A model is defined to be an “aligned” model if the model fulfills three criteria of helpful, honest, and harmless or “HHH” [95].
Researchers employ reinforcement learning with human feedback (RLHF) [96] for model alignment. In RLHF, a fine-tuned model on demonstrations is further trained with reward modeling (RM) and reinforcement learning (RL), shown in Figure 6. Below we briefly discuss RM and RL pipelines in RLHF.
Reward modeling: trains a model to rank generated responses according to human preferences using a classification objective. To train the classifier humans annotate LLMs generated responses based on HHH criteria.
Reinforcement learning: in combination with the reward model is used for alignment in the next stage. The previously trained reward model ranks LLM-generated responses into preferred vs. dispreferred, which is used to align the model with proximal policy optimization (PPO). This process repeats iteratively until convergence.
II-L3 Prompting/Utilization
Prompting is a method to query trained LLMs for generating responses, as illustrated in Figure 6. LLMs can be prompted in various prompt setups, where they can be adapted to the instructions without fine-tuning and in other cases with fine-tuning on data containing different prompt styles [16, 97, 98]. A good guide on prompt engineering is available at [32]. Below, we will discuss various widely used prompt setups.
 Zero-Shot Prompting: LLMs are zero-shot learners and capable of answering queries never seen before. This style of prompting requires LLMs to answer user questions without seeing any examples in the prompt.
 In-context Learning: Also known as few-shot learning, here, multiple input-output demonstration pairs are shown to the model to generate the desired response. This adaptation style is also called few-shot learning. A discussion on formatting in-context learning (ICL) templates is available in [50, 46, 18, 16].
  Reasoning in LLMs: LLMs are zero-shot reasoners and can be provoked to generate answers to logical problems, task planning, critical thinking, etc. with reasoning. Generating reasons is possible only by using different prompting styles, whereas to improve LLMs further on reasoning tasks many methods [16, 93] train them on reasoning datasets. We discuss various prompting techniques for reasoning below.
Chain-of-Thought (CoT): A special case of prompting where demonstrations contain reasoning information aggregated with inputs and outputs so that the model generates outcomes with step-by-step reasoning. More details on CoT prompts are available in [51, 99, 97].
Self-Consistency: Improves CoT performance by generating multiple responses and selecting the most frequent answer [100].
Tree-of-Thought (ToT): Explores multiple reasoning paths with possibilities to look ahead and backtrack for problem-solving [101].
Single-Turn Instructions: In this prompting setup, LLMs are queried only once with all the relevant information in the prompt. LLMs generate responses by understanding the context either in a zero-shot or few-shot setting.
Multi-Turn Instructions: Solving a complex task requires multiple interactions with LLMs, where feedback and responses from the other tools are given as input to the LLM for the next rounds. This style of using LLMs in the loop is common in autonomous agents.
III Large Language Models
This section reviews LLMs, briefly describing their architectures, training objectives, pipelines, datasets, and fine-tuning details.
III-A Pre-Trained LLMs
Here, we provide summaries of various well-known pre-trained LLMs with significant discoveries, changing the course of research and development in NLP. These LLMs have considerably improved the performance in NLU and NLG domains, and are widely fine-tuned for downstream tasks.
III-A1 General Purpose
T5 [10]
An encoder-decoder model employing a unified text-to-text training for all NLP problems, shown in Figure 7. T5 places layer normalization outside the residual path in a conventional transformer model [62]. It uses masked language modeling as a pre-training objective where spans (consecutive tokens) are replaced with a single mask instead of separate masks for each token. This type of masking speeds up the training as it produces shorter sequences. After pre-training, the model is fine-tuned using adapter layers [102] for downstream tasks.
Figure 7: Unified text-to-text training example, source image from [10].
GPT-3 [6]
The GPT-3 architecture is same as the GPT-2 [5] but with dense and sparse attention in transformer layers similar to the Sparse Transformer [63]. It shows that large models can train on larger batch sizes with a lower learning rate; in order to decide the batch size during training, GPT-3 uses the gradient noise scale as in  [103]. Overall, GPT-3 increases model parameters to 175B showing that the performance of large language models improves with the scale and is competitive with the fine-tuned models.
mT5 [11]
A multilingual T5 model [10] trained on the mC4 dataset with 101 languages. The dataset is extracted from the public common crawl scrape. The model uses a larger vocabulary size of 250,000 to cover multiple languages. To avoid over-fitting or under-fitting for a language, mT5 employs a data sampling procedure to select samples from all languages. The paper suggests using a small amount of pre-training datasets, including all languages when fine-tuning for a task using English language data. This allows the model to generate correct non-English outputs.
PanGu-$\alpha$ [104]
An autoregressive model that has a query layer at the end of standard transformer layers, example shown in Figure 8, with aim to predict next token. Its structure is similar to the transformer layer but with an additional embedding for the next position in the attention mechanism, given in Eq. 7.
$a=p_{n}W_{h}^{q}W_{h}^{k}TH_{L}^{T}$ (7)
Figure 8: The image is the article of [104], showing an example of PanGu-$\alpha$ architecture.
CPM-2 [12]
Cost-efficient Pre-trained language Models (CPM-2) pre-trains bilingual (English and Chinese) 11B and 198B mixture-of-experts (MoE) models on the WuDaoCorpus [105] dataset. The tokenization process removes “_” white space tokens in the sentencepiece tokenizer. The models are trained with knowledge inheritance, starting with only the Chinese language in the first stage and then adding English and Chinese data. This trained model gets duplicated multiple times to initialize the 198B MoE model. Moreover, to use the model for downstream tasks, CPM-2 experimented with both complete fine-tuning and prompt fine-tuning as in [106] where only prompt-related parameters are updated by inserting prompts at various positions, front, middle, and back. CPM-2 also proposes INFMOE, a memory-efficient framework with a strategy to dynamically offload parameters to the CPU for inference at a 100B scale. It overlaps data movement with inference computation for lower inference time.
ERNIE 3.0 [107]
ERNIE 3.0 takes inspiration from multi-task learning to build a modular architecture using Transformer-XL [108] as the backbone. The universal representation module is shared by all the tasks, which serve as the basic block for task-specific representation modules, which are all trained jointly for natural language understanding, natural language generation, and knowledge extraction. This LLM is primarily focused on the Chinese language, claims to train on the largest Chinese text corpora for LLM training, and achieved state-of-the-art in 54 Chinese NLP tasks.
Jurassic-1 [109]
A pair of auto-regressive language models, including a 7B-parameter J1-Large model and a 178B-parameter J1-Jumbo model. The training vocabulary of Jurassic-1 comprise word pieces, complete words, and multi-word expressions without any word boundaries, where possible out-of-vocabulary instances are interpreted as Unicode bytes. Compared to the GPT-3 counterparts, the Jurassic-1 models apply a more balanced depth-to-width self-attention architecture [110] and an improved tokenizer for a faster prediction based on broader resources, achieving a comparable performance in zero-shot learning tasks and a superior performance in few-shot learning tasks given the ability to feed more examples as a prompt.
HyperCLOVA [111]
A Korean language model with GPT-3 architecture.
Yuan 1.0 [112]
Trained on a Chinese corpus with 5TB of high-quality text collected from the Internet. A Massive Data Filtering System (MDFS) built on Spark is developed to process the raw data via coarse and fine filtering techniques. To speed up the training of Yuan 1.0 with the aim of saving energy expenses and carbon emissions, various factors that improve the performance of distributed training are incorporated in architecture and training like increasing the number of hidden size improves pipeline and tensor parallelism performance, larger micro batches improve pipeline parallelism performance, and higher global batch size improve data parallelism performance. In practice, the Yuan 1.0 model performs well on text classification, Winograd Schema, natural language inference, and reading comprehension tasks.
Gopher [113]
The Gopher family of models ranges from 44M to 280B parameters in size to study the effect of scale on the LLMs performance. The 280B model beats GPT-3 [6], Jurrasic-1 [109], MT-NLG [114], and others on 81% of the evaluated tasks.
ERNIE 3.0 TITAN [35]
ERNIE 3.0 Titan extends ERNIE 3.0 by training a larger model with 26x the number of parameters of the latter. This bigger model outperformed other state-of-the-art models in 68 NLP tasks. LLMs produce text with incorrect facts. In order to have control of the generated text with factual consistency, ERNIE 3.0 Titan adds another task, Credible and Controllable Generations, to its multi-task learning setup. It introduces additional self-supervised adversarial and controllable language modeling losses to the pre-training step, which enables ERNIE 3.0 Titan to beat other LLMs in their manually selected Factual QA task set evaluations.
GPT-NeoX-20B [115]
An auto-regressive model that largely follows GPT-3 with a few deviations in architecture design, trained on the Pile dataset without any data deduplication. GPT-NeoX has parallel attention and feed-forward layers in a transformer block, given in Eq. 8, that increases throughput by 15%. It uses rotary positional embedding [66], applying it to only 25% of embedding vector dimension as in [116]. This reduces the computation without performance degradation. Opposite to GPT-3, which uses dense and sparse layers, GPT-NeoX-20B uses only dense layers. The hyperparameter tuning at this scale is difficult; therefore, the model chooses hyperparameters from the method [6] and interpolates values between 13B and 175B models for the 20B model. The model training is distributed among GPUs using both tensor and pipeline parallelism.
$x+Attn(LN_{1}(x))+FF(LN_{2}(x))$ (8)
OPT [14]
It is a clone of GPT-3, developed with the intention to open-source a model that replicates GPT-3 performance. Training of OPT employs dynamic loss scaling  [117] and restarts from an earlier checkpoint with a lower learning rate whenever loss divergence is observed. Overall, the performance of OPT-175B models is comparable to the GPT3-175B model.
BLOOM [13]
A causal decoder model trained on ROOTS corpus with the aim of open-sourcing an LLM. The architecture of BLOOM is shown in Figure 9, with differences like ALiBi positional embedding, an additional normalization layer after the embedding layer as suggested by the bitsandbytes [^1] https://github.com/TimDettmers/bitsandbytes library. These changes stabilize training with improved downstream performance.
GLaM [118]
Generalist Language Model (GLaM) represents a family of language models using a sparsely activated decoder-only mixture-of-experts (MoE) structure [119, 120]. To gain more model capacity while reducing computation, the experts are sparsely activated where only the best two experts are used to process each input token. The largest GLaM model, GLaM (64B/64E), is about 7$\times$ larger than GPT-3 [6], while only a part of the parameters is activated per input token. The largest GLaM (64B/64E) model achieves better overall results as compared to GPT-3 while consuming only one-third of GPT-3’s training energy.
Figure 9: The BLOOM architecture example sourced from [13].
MT-NLG [114]
A 530B causal decoder based on GPT-2 architecture that is roughly 3$\times$ GPT-3 model parameters. MT-NLG is trained on filtered high-quality data collected from various public datasets and blends various types of datasets in a single batch, which beats GPT-3 on a number of evaluations.
Chinchilla [121]
A causal decoder trained on the same dataset as the Gopher [113] but with a little different data sampling distribution (sampled from MassiveText). The model architecture is similar to the one used for Gopher, with the exception of AdamW optimizer instead of Adam. Chinchilla identifies the relationship that model size should be doubled for every doubling of training tokens. Over 400 language models ranging from 70 million to over 16 billion parameters on 5 to 500 billion tokens are trained to get the estimates for compute-optimal training under a given budget. The authors train a 70B model with the same compute budget as Gopher (280B) but with 4 times more data. It outperforms Gopher [113], GPT-3 [6], and others on various downstream tasks, after fine-tuning.
AlexaTM [122]
An encoder-decoder model, where encoder weights and decoder embeddings are initialized with a pre-trained encoder to speedup training. The encoder stays frozen for initial 100k steps and later unfreezed for end-to-end training. The model is trained on a combination of denoising and causal language modeling (CLM) objectives, concatenating $[CLM]$ token at the beginning for mode switiching. During training, the CLM task is applied for 20% of the time, which improves the in-context learning performance.
PaLM [15]
A causal decoder with parallel attention and feed-forward layers similar to Eq. 8, speeding up training 15 times faster. Additional changes to the conventional transformer model include SwiGLU activation, RoPE embeddings, multi-query attention that saves computation cost during decoding, and shared input-output embeddings. During training, loss spiking was observed, and to fix it, model training was restarted from a 100 steps earlier checkpoint by skipping 200-500 batches around the spike. Moreover, the model was found to memorize around 2.4% of the training data at the 540B model scale, whereas this number was lower for smaller models.
  PaLM-2 [123]: A smaller multi-lingual variant of PaLM, trained for larger iterations on a better quality dataset. The PaLM-2 shows significant improvements over PaLM, while reducing training and inference costs due to its smaller size. To lessen toxicity and memorization, it appends special tokens with a fraction of pre-training data, which shows reduction in generating harmful responses.
U-PaLM [124]
This method trains PaLM for 0.1% additional compute with UL2 (also named as UL2Restore) objective [89] using the same dataset and outperforms baseline significantly on various NLP tasks, including zero-shot, few-shot, commonsense reasoning, CoT, etc. Training with UL2R involves converting a causal decoder PaLM to a non-causal decoder PaLM and employing 50% sequential denoising, 25% regular denoising, and 25% extreme denoising loss functions.
UL2 [89]
An encoder-decoder architecture trained using a mixture of denoisers (MoD) objectives. Denoisers include 1) R-Denoiser: a regular span masking, 2) S-Denoiser: which corrupts consecutive tokens of a large sequence and 3) X-Denoiser: which corrupts a large number of tokens randomly. During pre-training, UL2 includes a denoiser token from ${R,S,X}$ to represent a denoising setup. It helps improve fine-tuning performance for downstream tasks that bind the task to one of the upstream training modes. This MoD style of training outperforms the T5 model on many benchmarks.
GLM-130B [33]
GLM-130B is a bilingual (English and Chinese) model trained using an auto-regressive mask infilling pre-training objective similar to the GLM [125]. This training style makes the model bidirectional as compared to GPT-3, which is unidirectional. Opposite to the GLM, the training of GLM-130B includes a small amount of multi-task instruction pre-training data (5% of the total data) along with the self-supervised mask infilling. To stabilize the training, it applies embedding layer gradient shrink.
LLaMA [126, 21]
A set of decoder-only language models varying from 7B to 70B parameters. LLaMA models series is the most famous among the community for parameter-efficient and instruction tuning.
  LLaMA-1 [126]: Implements efficient causal attention [127] by not storing and computing masked attention weights and key/query scores. Another optimization is reducing number of activations recomputed in backward pass, as in  [128].
  LLaMA-2 [21]: This work is more focused towards fine-tuning a safer and better LLaMA-2-Chat model for dialogue generation. The pre-trained model has 40% more training data with a larger context length and grouped-query attention.
PanGu-$\Sigma$ [129]
An autoregressive model with parameters copied from PanGu-$\alpha$ and extended to a trillion scale with Random Routed Experts (RRE), the architectural diagram is shown in Figure 10. RRE is similar to the MoE architecture, with distinctions at the second level, where tokens are randomly routed to experts in a domain instead of using a learnable gating method. The model has bottom layers densely activated and shared across all domains, whereas top layers are sparsely activated according to the domain. This training style allows extracting task-specific models and reduces catastrophic forgetting effects in case of continual learning.
III-A2 Coding
CodeGen [130]
CodeGen has similar architecture to the PaLM [15], i.e., parallel attention, MLP layers, and RoPE embeddings. The model is trained on both natural language and programming language data sequentially (trained on the first dataset, then the second and so on) on the following datasets 1) PILE, 2) BIGQUERY and 3) BIGPYTHON. CodeGen proposed a multi-step approach to synthesizing code. The purpose is to simplify the generation of long sequences where the previous prompt and generated code are given as input with the next prompt to generate the next code sequence. CodeGen opensource a Multi-Turn Programming Benchmark (MTPB) to evaluate multi-step program synthesis.
Codex [131]
This LLM is trained on a subset of public Python Github repositories to generate code from docstrings. Computer programming is an iterative process where the programs are often debugged and updated before fulfilling the requirements. Similarly to this, Codex generates 100 versions of a program by repetitive sampling for a given description, which produces a working solution for 77.5% of the problems passing unit tests. Its powerful version powers Github Copilot [^2] https://github.com/features/copilot.
AlphaCode [132]
A set of large language models, ranging from 300M to 41B parameters, designed for competition-level code generation tasks. It uses the multi-query attention [133] to reduce memory and cache costs. Since competitive programming problems highly require deep reasoning and an understanding of complex natural language algorithms, the AlphaCode models are pre-trained on filtered GitHub code in popular languages and then fine-tuned on a new competitive programming dataset named CodeContests. The CodeContests dataset mainly contains problems, solutions, and test cases collected from the Codeforces platform [^3] https://codeforces.com/. The pre-training employs standard language modeling objectives, while GOLD [134] with tempering [135] serves as the training objective for the fine-tuning on CodeContests data. To evaluate the performance of AlphaCode, simulated programming competitions are hosted on the Codeforces platform: overall, AlphaCode ranks at the top 54.3% among over 5000 competitors, where its Codeforces rating is within the top 28% of recently participated users.
CodeT5+ [34]
CodeT5+ is based on CodeT5 [136], with shallow encoder and deep decoder, trained in multiple stages initially unimodal data (code) and later bimodal data (text-code pairs). Each training stage has different training objectives and activates different model blocks encoder, decoder, or both according to the task. The unimodal pre-training includes span denoising and CLM objectives, whereas bimodal pre-training objectives contain contrastive learning, matching, and CLM for text-code pairs. CodeT5+ adds special tokens with the text to enable task modes, for example, $[CLS]$ for contrastive loss, $[Match]$ for text-code matching, etc.
StarCoder [137]
A decoder-only model with SantaCoder architecture, employing Flash attention to scale up the context length to 8k. The StarCoder trains an encoder to filter names, emails, and other personal data from the training data. Its fine-tuned variant outperforms PaLM, LLaMA, and LAMDA on HumanEval and MBPP benchmarks.
III-A3 Scientific Knowledge
Galactica [138]
A large curated corpus of human scientific knowledge with 48 million papers, textbooks, lecture notes, millions of compounds and proteins, scientific websites, encyclopedias, and more are trained using metaseq library3, which is built on PyTorch and fairscale [139]. The model wraps reasoning datasets with $<work>$ token to provide step-by-step reasoning context to the model, which has been shown to improve the performance on reasoning tasks.
III-A4 Dialog
LaMDA [140]
A decoder-only model pre-trained on public dialog data, public dialog utterances, and public web documents, where more than 90% of the pre-training data is in English. LaMDA is trained with the objective of producing responses that exhibit high levels of quality, safety, and groundedness. To achieve this, discriminative and generative fine-tuning techniques are incorporated to enhance the model’s safety and quality aspects. As a result, the LaMDA models can be utilized as a general language model performing various tasks.
III-A5 Finance
BloombergGPT [141]
A non-causal decoder model trained using both financial ("FINPILE" from the Bloomberg archive) and general-purpose datasets. The model’s architecture is similar to the BLOOM [13] and OPT [14]. It allocates 50B parameters to different blocks of the model using the approach [142]. For effective training, BloombergGPT packs documents together with $<|endoftext|>$ to use maximum sequence length, use warmup batch size starting from 1024 to 2048, and manually reduces the learning rate multiple times during the training.
Xuan Yuan 2.0 [143]
A Chinese financial chat model with BLOOM’s [13] architecture trained on a combination of general purpose, financial, general purpose instructions, and financial institutions datasets. Xuan Yuan 2.0 combined the pre-training and fine-tuning stages to avoid catastrophic forgetting.
Figure 10: This example illustrates the PanGu-$\sum$ architecture, as depicted in the image sourced from [129].
TABLE I: Noteworthy findings and insights from pre-trained Large Language Model.
Models Findings & Insights
T5
Encoder and decoder with shared parameters perform equivalently when parameters are not shared Fine-tuning model layers (adapter layers) work better than the conventional way of training on only classification layers
GPT-3
Few-shot performance of LLMs is better than the zero-shot, suggesting that LLMs are meta-learners
mT5
Large multi-lingual models perform equivalently to single language models on downstream tasks. However, smaller multi-lingual models perform worse
PanGu-$\alpha$
LLMs are good at a few shot capabilities
CPM-2
Prompt fine-tuning requires updating very few parameters while achieving performance comparable to full model fine-tuning Prompt fine-tuning takes more time to converge as compared to full model fine-tuning Inserting prompt tokens in-between sentences can allow the model to understand relations between sentences and long sequences In an analysis, CPM-2 finds that prompts work as a provider (additional context) and aggregator (aggregate information with the input text) for the model
Codex
This LLM focuses on code evaluations and introduces a novel way of selecting the best code samples. The results indicate it is possible to accurately select code samples using heuristic ranking in lieu of a detailed evaluation of each sample, which may not be feasible or feasible in some situations.
ERNIE 3.0
ERNIE 3.0 shows that a modular LLM architecture with a universal representation module and task-specific representation module helps in finetuning phase. Optimizing the parameters of a task-specific representation network during the fine-tuning phase is an efficient way to take advantage of the powerful pretrained model.
Jurassic-1
The performance of an LLM is highly related to the network size. To improve runtime performance, more operations can be performed in parallel (width) rather than sequentially (depth). To efficiently represent and fit more text in the same context length, the model uses a larger vocabulary to train a SentencePiece tokenizer without restricting it to word boundaries. This tokenizer improvement can further benefit few-shot learning tasks.
HyperCLOVA
By employing prompt-based tuning, the performances of models can be improved, often surpassing those of state-of-the-art models when the backward gradients of inputs are accessible.
Yuan 1.0
The model architecture that excels in pre-training and fine-tuning cases may exhibit contrasting behavior in zero-shot and few-shot learning.
Gopher
Relative encodings enable models to be evaluated for longer sequences than those on which it was trained.
ERNIE 3.0 Titan
This LLM builds on top of ERNIE 3.0 and add a self-supervised adversarial loss to distinguish whether a text is generated or the original one. This distinction ability between real and generate text improves the LLM’s performance as compared to ERNIE 3.0.
GPT-NeoX-20B
Parallel attention + FF layers speed-up training 15% with the same performance as with cascaded layers Initializing feed-forward output layers before residuals with scheme in [144] avoids activations from growing with increasing depth and width Training on Pile outperforms GPT-3 on five-shot
OPT
Restart training from an earlier checkpoint with a lower learning rate if loss diverges Model is prone to generate repetitive text and stuck in a loop
BLOOM
None
Galactica
Galactica’s performance has continued to improve across validation set, in-domain, and out-of-domain benchmarks, even with multiple repetitions of the corpus, which is superior to existing research on LLMs. A working memory token approach can achieve strong performance over existing methods on mathematical MMLU and MATH benchmarks. It sets a new state-of-the-art on several downstream tasks such as PubMedQA (77.6%) and MedMCQA dev (52.9%).
GLaM
The feed-forward component of each Transformer layer can be replaced with a mixture-of-experts (MoE) module consisting of a set of independent feed-forward networks (i.e., the ‘experts’). By sparsely activating these experts, the model capacity can be maintained while much computation is saved. By leveraging sparsity, we can make significant strides toward developing high-quality NLP models while simultaneously reducing energy consumption. Consequently, MoE emerges as a robust candidate for future scaling endeavors. The model trained on filtered data shows consistently better performances on both NLG and NLU tasks, where the effect of filtering is more significant on the former tasks. Filtered pretraining corpora plays a crucial role in the generation capability of LLMs, especially for the downstream tasks. The scaling of GLaM MoE models can be achieved by increasing the size or number of experts in the MoE layer. Given a fixed budget of computation, more experts contribute to better predictions.
LaMDA
The model can be fine-tuned to learn to call different external information resources and tools.
MT-NLG
None.
AlphaCode
For higher effectiveness and efficiency, a transformer model can be asymmetrically constructed with a shallower encoder and a deeper decoder. To achieve better performances, it is necessary to employ strategies such as massively scaling up sampling, followed by the filtering and clustering of samples into a compact set. The utilization of novel sampling-efficient transformer architectures designed to facilitate large-scale sampling is crucial. Simplifying problem descriptions can effectively improve the model’s performance.
Table Continued on Next Page
Models Findings & Insights
Chinchilla
The experiments that culminated in the development of Chinchilla determined that for optimal computation during training, the model size and the number of training tokens should be scaled proportionately: for each doubling of the model size, the number of training tokens should be doubled as well.
PaLM
English-centric models produce better translations when translating to English as compared to non-English Generalized models can have equivalent performance for language translation to specialized small models Larger models have a higher percentage of training data memorization Performance has not yet saturated even at 540B scale, which means larger models are likely to perform better
AlexaTM
Compared to commonly used Decoder-only Transformer models, seq2seq architecture is more suitable for training generative LLMs given stronger bidirectional attention to the context. An extra Causal Language Modeling (CLM) task can be added to benefit the model with a more efficient in-context learning, especially for few-shot learning tasks. The key to training powerful seq2seq-based LLMs lies in mixed pre-training, rather than additional multitask training. Placing layernorms at the beginning of each transformer layer can improve the training stability of large models.
U-PaLM
Training with a mixture of denoisers outperforms PaLM when trained further for a few more FLOPs Training with a mixture of denoisers improves the infilling ability and open-ended text generation diversity
UL2
Mode switching training enables better performance on downstream tasks CoT prompting outperforms standard prompting for UL2
GLM-130B
Pre-training data with a small proportion of multi-task instruction data improves the overall model performance
CodeGen
Multi-step prompting for code synthesis leads to a better user intent understanding and code generation
LLaMA
LLaMA is open-source and can be fine-tuned or continually pre-trained to develop new models or instruction-based tools. A few optimizations are proposed to improve the training efficiency of LLaMA, such as efficient implementation of multi-head self-attention and a reduced amount of activations during back-propagation. Training exclusively on public data can also achieve state-of-the-art performance. A constant performance improvement is gained when scaling the model. Smaller models can also realize good performances using more training data and time.
PanGu-$\Sigma$
Sparse models provide the benefits of large models at a lower computation cost Randomly Routed Experts reduces catastrophic forgetting effects which in turn is essential for continual learning Randomly Routed Experts allow extracting a domain-specific sub-model in deployment which is cost-efficient while maintaining a performance similar to the original
BloombergGPT
Pre-training with general-purpose and task-specific data improves task performance without hurting other model capabilities
XuanYuan 2.0
Combining pre-training and fine-tuning stages in single training avoids catastrophic forgetting
CodeT5+
Causal LM is crucial for a model’s generation capability in encoder-decoder architectures Multiple training objectives like span corruption, Causal LM, matching, etc complement each other for better performance
StarCoder
HHH prompt by Anthropic allows the model to follow instructions without fine-tuning
LLaMA-2
Model trained on unfiltered data is more toxic but may perform better on downstream tasks after fine-tuning Model trained on unfiltered data requires fewer samples for safety alignment
PaLM-2
Data quality is important to train better models Model and data size should be scaled with 1:1 proportions Smaller models trained for larger iterations outperform larger models
TABLE II: Key insights and findings from the study of instruction-tuned Large Language Models.
Models Findings & Insights
T0
Multi-task prompting enables zero-shot generalization and outperforms baselines Even a single prompt per dataset task is enough to improve performance
WebGPT
The answer quality of LLMs can be further improved with human feedback. To aid the model in effectively filtering and utilizing relevant information, human labelers play a crucial role in answering questions regarding the usefulness of the retrieved documents. Interacting a fine-tuned language model with a text-based web-browsing environment can improve end-to-end retrieval and synthesis via imitation learning and reinforcement learning. Generating answers with references can make labelers easily judge the factual accuracy of answers.
Tk-INSTRUCT
Instruction tuning leads to a stronger generalization of unseen tasks More tasks improve generalization whereas only increasing task instances does not help Supervised trained models are better than generalized models Models pre-trained with instructions and examples perform well for different types of inputs
mT0 and BLOOMZ
Instruction tuning enables zero-shot generalization to the tasks never seen before Multi-lingual training leads to even better zero-shot generalization for both English and non-English Training on machine-translated prompts improves performance for held-out tasks with non-English prompts English only fine-tuning on multilingual pre-trained language model is enough to generalize to other pre-trained language tasks
OPT-IML
Task size sampling to create a batch with most of the task examples is important for better performance Only example proportional sampling is not enough, training datasets/benchmarks should also be proportional for better generalization/performance Fully held-out and partially supervised tasks performance improves by scaling tasks or categories whereas fully supervised tasks have no effect Including small amounts i.e. 5% of pretraining data during fine-tuning is effective Only 1% reasoning data improves the performance, adding more deteriorates performance Adding dialogue data makes the performance worse
Flan
Finetuning with CoT improves performance on held-out tasks Fine-tuning along with CoT data improves reasoning abilities CoT tuning improves zero-shot reasoning Performance improves with more tasks Instruction fine-tuning improves usability which otherwise is challenging for pre-trained models Improving the model’s performance with instruction tuning is compute-efficient Multitask prompting enables zero-shot generalization abilities in LLM
Sparrow
The judgments of labelers and the alignments with defined rules can help the model generate better responses. Good dialogue goals can be broken down into detailed natural language rules for the agent and the raters. The combination of reinforcement learning (RL) with reranking yields optimal performance in terms of preference win rates and resilience against adversarial probing.
WizardCoder
Fine-tuning with re-written instruction-tuning data into a complex set improves the performance significantly
LLaMA-2-Chat
Model learns to write safe responses with fine-tuning on safe demonstrations, while additional RLHF step further improves model safety and make it less prone to jailbreak attacks
LIMA
Less high quality data is enough for fine-tuned model generalization
III-B Fine-Tuned LLMs
Pre-trained LLMs have excellent generalization abilities to unseen tasks. However, because they are generally trained with the objective of next token prediction, LLMs have limited capacity to follow user intent and are prone to generate unethical, toxic or inaccurate responses [20]. For their effective utilization, LLMs are fine-tuned to follow instructions [16, 17, 93] and generate safe responses [20], which also results in increasing zero-shot, few-shot, and cross-task generalization [93, 16, 18], with minimal compute increment, e.g., 0.2% of the total pre-training for PaLM 540B [16].
We review various fine-tuned LLMs and strategies for effective fine-tuning in this section.
III-B1 Instruction-Tuning with Manually Created Datasets
Numerous hand-crafted instruction-tuning datasets with different design choices are proposed in the literature to instruction-tune LLMs. The performance of fine-tuned LLMs depends on multiple factors, such as dataset, instruction diversity, prompting templates, model size, and training objectives. Keeping this in view, diverse fine-tuned models have emerged in the literature using manually created datasets.
The models T0 [17] and mT0 (multi-lingual) [145] employ templates to convert existing datasets into prompt datasets. They have shown improvements in generalization to zero-shot and held-out tasks. Tk-Instruct [18] fine-tuned the T5 model with in-context instructions to study generalization on unseen tasks when given in-context instructions during test time. The model outperformed Instruct-GPT, despite being smaller in size, i.e., 11B parameters as compared to 175B of GPT-3.
Increasing Tasks and Prompt Setups: Zero-shot and few-shot performance improves significantly by expanding task collection and prompt styles. OPT-IML [93] and Flan [16] curated larger 2k and 1.8k task datasets, respectively. While increasing task size alone is not enough, OPT-IML and Flan add more prompting setups in their datasets, zero-shot, few-shot, and CoT. In continuation, CoT Collection [97] fine-tunes Flan-T5 further on 1.88M CoT samples. Another method [98] uses symbolic tasks with tasks in T0, Flan, etc.
Figure 11: An example image shows an instance of the Flan training paradigm, taken from [16].
III-B2 Instruction-Tuning with LLMs Generated Datasets
Generating an instruction-tuning dataset requires carefully writing instructions and input-output pairs, which are often written by humans, smaller in size, and less diverse. To overcome this, self-instruct [19] proposed an approach to prompt available LLMs to generate instruction-tuning datasets. Self-instruct outperformed models trained on manually created dataset SUPER-NATURALINSTRUCTIONS (a dataset with 1600+ tasks) [18] by 33%. It starts with a seed of 175 tasks, 1 instruction, and 1 sample per task and iteratively generates new instructions (52k) and instances (82k input-output pairs) using GPT-3 [6]. Contrary to this, Dynosaur [146] uses the meta-data of datasets on Huggingface to prompt LLMs to generate multiple task instruction-tuning datasets.
LLaMA Tuned: Various models in literature instruction-tune LLaMA [147] with GPT-3 [6] or GPT-4 [148] generated datasets. Among these, Alpaca [149], Vicuna [150], and LLaMA-GPT-4 [151] are a few general-purpose fine-tuned models, where Alpaca is trained on 52k samples from text-davinci-003, Vicuna on 70k samples from ShareGPT.com, and LLaMA-GPT-4 by re-creating Alpaca instructions from GPT-4. Goat [152] fine-tunes LLaMA for arithmetic tasks (1 million samples) by generating data from ChatGPT and outperforms GPT-4, PaLM, BLOOM, OPT, etc, attributing its success to the LLaMA’s consistent tokenization of numbers. HuaTuo [153] is a medical knowledge model, fine-tuned with a generated QA dataset of 8k instructions.
Complex Instructions: Evol-Instruct [154, 155] prompts LLMs to convert given instructions into a more complex set. The instructions are iteratively evolved with re-writing instructions in complex wording and creating new instructions. With this style of automated instruction generation, WizardLM [154] (fine-tuned LLaMA on 250k instructions), outperforms Vicuna and Alpaca, and WizardCoder [155] (fine-tuned StarCoder) beats Claude-Plus, Bard, and others.
III-B3 Aligning with Human Preferences
Incorporating human preferences into LLMs presents a significant advantage in mitigating undesirable behaviors and ensuring accurate outputs. The initial work on alignment, such as InstructGPT [20] aligns GPT-3 using a 3-step approach, instruction-tuning, reward modeling, and fine-tuning with reinforcement learning (RL). The supervised fine-tuned GPT-3 on demonstrations is queried to generate responses, which human labelers rank according to human values, and a reward model is trained on the ranked data. Lastly, the GPT-3 is trained with proximal policy optimization (PPO) using rewards on the generated data from the reward model. LLaMA 2-Chat [21] improves alignment by dividing reward modeling into helpfulness and safety rewards and using rejection sampling in addition to PPO. The initial four versions of LLaMA 2-Chat are fine-tuned with rejection sampling and then with PPO on top of rejection sampling.
  Aligning with Supported Evidence: This style of alignment allows the model to generate responses with proofs and facts, reduces hallucination, and assists humans more effectively, which increases trust in the model’s output. Similar to the RLHF training style, a reward model is trained to rank generated responses containing web citations in answers to questions, which is later used to train the model, as in GopherCite [156], WebGPT [157], and Sparrow [158]. The ranking model in Sparrow [158] is divided into two branches, preference reward and rule reward, where human annotators adversarial probe the model to break a rule. These two rewards together rank a response to train with RL.
  Aligning Directly with SFT: The PPO in the RLHF pipeline is complex, memory-intensive, and unstable, requiring multiple models, reward, value, policy, and reference models. Avoiding this sophisticated alignment pipeline is possible by incorporating minimal changes in the supervised fine-tuning (SFT) pipeline as in [159, 160, 161], with better or comparable performance to PPO. Direct preference optimization (DPO) [159] trains a model directly on the human-preferred responses to maximize the likelihood of preferred against unpreferred responses, with per-sample importance weight. Reward ranked fine-tuning RAFT [160] fine-tunes the model on ranked responses by the reward model. Preference ranking optimization (PRO) [162] and RRHF [161] penalize the model to rank responses with human preferences and supervised loss. On the other hand, chain-of-hindsight (CoH) [163] provides feedback to the model in language rather than reward, to learn good versus bad responses.
 Aligning with Synthetic Feedback: Aligning LLMs with human feedback is slow and costly. The literature suggests a semi-automated process to align LLMs by prompting LLMs to generate helpful, honest, and ethical responses to the queries, and fine-tuning using the newly created dataset. Constitutional AI [164] replaces human feedback in RLHF with AI, calling it RL from AI feedback (RLAIF). AlpacaFarm [165] designs prompts to imitate human feedback using LLMs APIs. Opposite to constitutional AI, AlpacaFarm injects noise in feedback to replicate human mistakes. Self-Align [94] prompts the LLM with ICL examples, instructing the LLM about what the response should contain to be considered useful and ethical. The same LLM is later fine-tuned with the new dataset.
  Aligning with Prompts: LLMs can be steered with prompts to generate desirable responses without training [166, 167]. The self-correction prompting in [167] concatenates instructions and CoT with questions, guiding the model to answer its instruction following strategy to ensure moral safety before the actual answer. This strategy is shown to reduce the harm in generated responses significantly.
  Red-Teaming/Jailbreaking/Adversarial Attacks: LLMs exhibit harmful behaviors, hallucinations, leaking personal information, and other shortcomings through adversarial probing. The models are susceptible to generating harmful responses even though they are aligned for safety [168, 169]. Red-teaming is a common approach to address illicit outputs, where the LLMs are prompted to generate harmful outputs [169, 170]. The dataset collected through red-teaming is used to fine-tune models for safety. While red-teaming largely relies on human annotators, another work [171] red-team LLMs to find prompts that lead to harmful outputs of other LLMs.
III-B4 Continue Pre-Training
Although fine-tuning boosts a model’s performance, it leads to catastrophic forgetting of previously learned information. Concatenating fine-tuning data with a few randomly selected pre-training samples in every iteration avoids network forgetting [172, 143]. This is also effective in adapting LLMs for cases where fine-tuning data is small and the original capacity is to be maintained. Prompt-based continued pre-training (PCP) [173] trains the model with text and instructions related to tasks and then finally instruction-tunes the model for downstream tasks.
III-B5 Sample Efficiency
While fine-tuning data is generally many-fold smaller than the pre-training data, it still has to be large enough for acceptable performance [16, 93, 18] and requires proportional computing resources. To study the effects on performance with less data, existing literature [174, 175] finds that the models trained on lesser data can outperform models trained with more data. In [174], 25% of the total downstream data is found enough for state-of-the-art performance. Selecting coreset-based 0.5% of the total instruction-tuning data improves the model performance by 2% in [175], as compared to the complete data tuning. Less is more for alignment (LIMA) [176] uses only 1000 carefully created demonstrations to fine-tune the model and has achieved comparable performance to GPT-4.
III-C Increasing Context Window
LLMs are trained with limited context windows due to expensive attention and high memory requirements. A model trained on limited sequence lengths fails to generalize to unseen lengths at inference time [177, 45]. Alternatively, LLMs with ALiBi [65] positional encodings can perform zero-shot length extrapolation. However, ALiBi has less expressive power [66] and inferior performance on multiple benchmarks [42], and many LLMs use RoPE positional embedding that is unable to perform zero-shot extrapolation. A larger context length has benefits such as a better understanding of longer documents, more samples in in-context learning, execution of bigger reasoning processes, etc. Expanding context length during fine-tuning is slow, inefficient, and computationally expensive [45]. Therefore, researchers employ various context window extrapolation techniques discussed below.
Position Interpolation: Rather than extrapolating, [45] shows that interpolating position encodings within the pre-trained context window are more effective. The work demonstrates that only 1000 steps of fine-tuning are enough to achieve better results on larger windows without performance loss compared to the original context size. Giraffe [42] uses power scaling in RoPE, and YaRN [43] proposed NTK-aware interpolation.
Efficient Attention Mechanism: Dense global attention is one of the major constraints in training larger context window LLMs. Using efficient attention variants, such as local, sparse, and dilated attention, reduces the computation cost significantly. LongT5 [44] proposes transient global attention (TGlobal), applying attention to local and global tokens (windowing token averaging). The model replaces attention in T5 [10] with TGlobal attention, pre-trains the model on 4098 sequence length, fine-tunes on larger window sizes, as large as 16k, and improves task performance with longer inputs. This shows the extrapolation ability of TGlobal attention with only fine-tuning. COLT5 [178] uses two branches, one with lightweight and the other with heavyweight attention and feed-forward layers. All tokens are processed from the lightweight branch, and only important tokens are routed to the heavyweight branch. LongNet [179] replaces standard attention with dilated attention, expanding sequence length to 1 billion tokens. LongLoRA [180] proposes shift-short attention, used during fine-tuning to reduce dense attention costs, while the model during inference can use dense attention and achieve similar performance as full attention fine-tuning.
Extrapolation without Training: LM-Infinite [177] and parallel context windows (PCW) [181] show length extrapolation is possible using pre-trained LLMs. LM-Infinite suggested $\Lambda$-shaped attention applied within the original context window limits. Likewise, PCW chunks larger inputs into the pre-trained context lengths and applies the same positional encodings to each chunk.
III-D Robotics
LLMs have been rapidly adopted across various domains in the scientific community due to their multipurpose capabilities [46]. In robotics research, the LLMs have very promising applications as well, such as enhancing human-robot interaction [28, 182, 183, 184], task planning [185, 186, 187], navigation [188, 189], and learning [190, 191]. They can enable robots to understand and generate natural language, aiding in instruction following, data annotation, and collaborative problem-solving. They can facilitate continuous learning by allowing robots to access and integrate information from a wide range of sources. This can help robots acquire new skills, adapt to changes, and refine their performance based on real-time data.
LLMs have also started assisting in simulating environments for testing and offer potential for innovative research in robotics, despite challenges like bias mitigation and integration complexity. The work in [192] focuses on personalizing robot household cleanup tasks. By combining language-based planning and perception with LLMs, such that having users provide object placement examples, which the LLM summarizes to generate generalized preferences, they show that robots can generalize user preferences from a few examples. An embodied LLM is introduced in [26], which employs a Transformer-based language model where sensor inputs are embedded alongside language tokens, enabling joint processing to enhance decision-making in real-world scenarios. The model is trained end-to-end for various embodied tasks, achieving positive transfer from diverse training across language and vision domains. LLMs have also been explored as zero-shot human models for enhancing human-robot interaction.
The study in [28] demonstrates that LLMs, trained on vast text data, can serve as effective human models for certain HRI tasks, achieving predictive performance comparable to specialized machine-learning models. However, limitations were identified, such as sensitivity to prompts and difficulties with spatial/numerical reasoning. In another study [193], the authors enable LLMs to reason over sources of natural language feedback, forming an “inner monologue” that enhances their ability to process and plan actions in robotic control scenarios. They combine LLMs with various forms of textual feedback, allowing the LLMs to incorporate conclusions into their decision-making process for improving the execution of user instructions in different domains, including simulated and real-world robotic tasks involving tabletop rearrangement and mobile manipulation. All of these studies employ LLMs as the core mechanism for assimilating everyday intuitive knowledge into the functionality of robotic systems.
Planning: LLMs are increasingly integral in robotics, particularly for strategic planning [185, 194, 195]. Their proficiency in processing and generating natural language is crucial for enhancing human-robot interaction and enabling robots to understand and execute complex tasks based on verbal instructions. LLMs also play a key role in task planning, a higher-level cognitive process involving the determination of sequential actions needed to achieve specific goals. This proficiency is crucial across a spectrum of applications, from autonomous manufacturing processes to household chores, where the ability to understand and execute multi-step instructions is of paramount significance.
Manipulation: In the area of manipulation [196, 197, 198, 199], LLMs enhance a robot’s dexterity and adaptability, excelling in tasks like object recognition, grasping, and collaboration. They analyze visual and spatial information to determine the most effective approach to interact with objects, proving invaluable in operations requiring precision and flexibility, such as surgical procedures or assembly line tasks. They also enable the integration of sensor inputs and linguistic cues in an embodied framework, enhancing decision-making in real-world scenarios. It enhances the model’s performance across various embodied tasks by allowing it to gather insights and generalize from diverse training data spanning language and vision domains.
Navigation: LLMs have revolutionized the navigation in robotics [200, 201, 202, 203], offering significant potential to enhance a robot’s ability to navigate complex environments with precision and adaptability. Motion planning [188], in particular, stands out as a critical domain where LLMs have shown remarkable promise, excelling in generating feasible paths and trajectories for robots, accounting for intricate environmental details. This ability proves particularly valuable in scenarios requiring precise and dynamically adaptable navigation, as observed in environments like warehouses, transport and healthcare facilities, and smart residences. LLMs have also played a key role in localization and mapping, which are foundational components for successful robot navigation. They empower robots to determine their precise position within an environment while concurrently constructing or updating a spatial representation of their surroundings. This capability is crucial for tasks demanding spatial awareness, including autonomous exploration, search and rescue missions, and the operations of mobile robots. They have also contributed significantly to the proficiency of collision-free navigation within the environment while accounting for obstacles and dynamic alterations, playing an important role in scenarios where robots are tasked with traversing predefined paths with accuracy and reliability, as seen in the operations of automated guided vehicles (AGVs) and delivery robots (e.g., SADRs – pedestrian sized robots that deliver items to customers without the involvement of a delivery person).
III-E Multimodal LLMs
Inspired by the success of LLMs in natural language processing applications, an increasing number of research works are now facilitating LLMs to perceive different modalities of information like image [204, 205, 206], video [207, 208, 209], audio [210, 209, 211], etc. Multimodal LLMs (MLLMs) present substantial benefits compared to standard LLMs that process only text. By incorporating information from various modalities, MLLMs can achieve a deeper understanding of context, leading to more intelligent responses infused with a variety of expressions. Importantly, MLLMs align closely with human perceptual experiences, leveraging the synergistic nature of our multisensory inputs to form a comprehensive understanding of the world [211, 26]. Coupled with a user-friendly interface, MLLMs can offer intuitive, flexible, and adaptable interactions, allowing users to engage with intelligent assistants through a spectrum of input methods. According to the ways of constructing models, current MLLMs can be generally divided into three streams: pre-training, fine-tuning, and prompting. In this section, we will discuss more details of these main streams, as well as the important application of MLLMs in visual reasoning.
Pre-training: This stream of MLLMs intends to support different modalities using unified end-to-end models. For instance, Flamingo [204] applies gated cross-attention to fuse vision and language modalities, which are collected from pre-trained and frozen visual encoder and LLM, respectively. Moreover, BLIP-2 [205] proposes a two-stage strategy to pre-train a Querying Transformer (Q-Former) for the alignment between vision and language modalities: in the first stage, vision-language representation learning is bootstrapped from a frozen visual encoder; and in the second stage, a frozen LLM bootstraps vision-to-language generative learning for zero-shot image-to-text generation. Similarly, MiniGPT-4 [212] also deploys pre-trained and frozen ViT [213], Q-Former and Vicuna LLM [150], while only a linear projection layer needs to be trained for vision and language modalities alignment.
Fine-tuning: Derived from instruction tuning [16] for NLP tasks [20, 16, 93], researchers are now fine-tuning pre-trained LLMs using multimodal instructions. Following this method, LLMs can be easily and effectively extended as multimodal chatbots [212, 206, 29] and multimodal task solvers [214, 30, 215]. The key issue of this stream of MLLMs is to collect multimodal instruction-following data for fine-tuning [216]. To address this issue, the solutions of benchmark adaptation [214, 217, 218], self-instruction [19, 31, 219], and hybrid composition [220, 215] are employed, respectively. To mitigate the gap between the original language modality and additional modalities, the learnable interface is introduced to connect different modalities from frozen pre-trained models. Particularly, the learnable interface is expected to work in a parameter-efficient tuning manner: e.g., LLaMA-Adapter [221] applies an efficient transformer-based adapter module for training, and LaVIN [220] dynamically learns the multimodal feature weights using a mixture-of-modality adapter. Different from the learnable interface, the expert models can directly convert multimodalities into language: e.g., VideoChat-Text [207] incorporates Whisper [222], a speech recognition expert model, to generate the captions of given videos for the understanding of following LLMs.
Prompting: Different from the fine-tuning technique that directly updates the model parameters given task-specific datasets, the prompting technique provides certain context, examples, or instructions to the model, fulfilling specialized tasks without changing the model parameters. Since prompting can significantly reduce the need for large-scale multimodal data, this technique is widely used to construct MLLMs. Particularly, to solve multimodal Chain of Thought (CoT) problems [99], LLMs are prompted to generate both the reasoning process and the answer given multimodal inputs [223]. On this front, different learning paradigms are exploited in practice: for example, Multimodal-CoT [223] involves two stages of rationale generation and answer inference, where the input of the second stage is a combination of the original input and the output of the first stage; and CoT-PT [224] applies both prompt tuning and specific visual bias to generate a chain of reasoning implicitly. In addition to CoT problems, LLMs can also be prompted with multimodal descriptions and tools, effectively dividing complex tasks into sub-tasks [225, 226].
Visual Reasoning Application: Recent visual reasoning systems [227, 228, 229, 230] tend to apply LLMs for better visual information analysis and visual-language integration. Different from previous works [231, 232] that rely on limited VQA datasets and small-scale neural networks, current LLM-aided methods offer benefits of stronger generalization ability, emergent ability, and interactivity [216]. To realize visual reasoning with the help of LLMs, prompting and fine-tuning techniques can also be utilized: for example, PointClip V2 [228] applies LLMs to generate 3D-specific prompts, which are encoded as textual features and then combined with visual features for 3D recognition; and GPT4Tools [31] employs LoRA [233] to fine-tune LLMs following tool-related instructions. Serving as a controller [230], decision maker [234], or semantics refiner [227, 235], LLMs significantly facilitates the progress of visual reasoning research.
III-F Augmented LLMs
LLMs are capable of learning from the examples concatenated with the input, known as context augmentation, in-context learning (ICL), or few-shot prompting. They show excellent generalization to unseen tasks with few-shot prompting, enabling LLMs to answer queries beyond the capacity acquired during training [6, 51]. These emergent abilities allow for adapting the model without fine-tuning - a costly process. Aside from this, hallucination, producing inaccurate, unsafe or factually incorrect responses, is common for LLMs, which is avoided by augmenting contextual data. While the user can provide in-context samples in the query [50, 32], here we specifically refer to the methods that access external storage programmatically, calling them augmented LLMs.
The literature suggests various external memory designs to augment LLMs, long-term [236, 237, 238, 239], short-term [240], symbolic [241], and non-symbolic [242, 243]. The memory can be maintained in different formats such as documents, vectors, or databases. A few systems maintain intermediate memory representations to retain information across multiple iterations [239, 237], while others extract important information from the datasets and save it in memory for recall [244]. The memory read and write operations are performed either with or without LLMs cooperation [237, 245, 239, 246], acting as a feedback signal in [240]. We discuss different types of augmented LLMs below.
Figure 12: A flow diagram of Retrieval Augmented LLMs. The retriever extracts a similar context to the input and forwards it to the LLM either in simple language or encoded through Fusion-in-Decoder (FiD). Depending on the task, retrieval and generation may repeat multiple times.
III-F1 Retrieval Augmented LLMs
LLMs may have limited memory and outdated information, leading to inaccurate responses. Retrieving relevant information from external up-to-date storage enables the LLMs to accurately answer with references and utilize more information. With retrieval augmentation, smaller models have been shown to perform at par with larger models. For instance, the 11B model can become competitive to 540B PaLM in [25] and 7.5B to 280B Gopher in [238]. Retrieval augmented language modeling (RALM) has two major components, shown in Figure 12, namely: 1) retriever and 2) language model. In RALM, the retriever plays a crucial role in driving LLM response, where incorrect information can steer LLMs to false behavior. This leads to the development of various methods to retrieve accurate information and fuse with the query for better performance.
Zero-Shot Retrieval Augmentation: This kind of augmentation keeps the original LLM architecture and weights unchanged and uses BM25 [247], nearest neighbors, or frozen pre-trained models like Bert [7] as a retriever. The retrieved information is provided as input to the model for response generation, shown to improve performance over LLMs without retrieval [243, 248]. In some scenarios, multiple retrieval iterations are required to complete the task. The output generated in the first iteration is forwarded to the retriever to fetch similar documents. Forward-looking active retrieval (FLARE) [242] initially generates the response and corrects the output by retrieving relevant documents if the response contains low-confidence tokens. Similarly, RepoCoder [249] fetches code snippets recursively for code completion.
Training with Retrieval Augmentation: To reduce failures in retrieval augmentation generation (RAG), researchers train or fine-tune retrievers and LLMs with a retrieval augmentation pipeline. We discuss the literature below based on their focus on the respective training processes of the pipeline.
Training LLM: Retrieval-enhanced transformer (RETRO) [238] shows pre-training smaller LLMs with RAG pipeline outperforms larger LLMs, such as GPT-3 trained without RAG. RETRO uses a 2-trillion token subset of MassiveText as a database. The retrieval pipeline divides the input query into subsets and retrieves relevant chunks from the database for each subset, encoded together with input intermediate representations for generating tokens. It uses cross-chunked attention to attend to previous chunks auto-regressively. A study on RETRO [250] shows models pre-trained without RAG but fine-tuned using RAG lack the performance gains obtained by pre-training with RAG.
Training Retriever: Quality of responses generated by LLMs is highly dependent on the in-context examples. Therefore, [251, 252, 253, 254] train retrievers to retrieve accurate few-shot samples while keeping the LLM frozen for generation. Retrieved samples are ranked to build ground-truth data to train retrievers with contrastive learning in [251, 253]. RoBERTa is trained for downstream tasks in [252] for ICL samples retrieval. REPLUG [254] trains the retriever with supervised signals from the frozen LLM-generated outputs.
Training Retriever and LLM: Further benefits are achieved by training both the retriever and the model in [25, 255, 256]. In this case, the error propagates back to the retriever, updating both the language model and the retriever. While masked language modeling (MLM) is a common pre-training objective [25, 256], retrieval pre-trained transformer (RPT) [255] used document chunk prediction as a pre-training objective for long text modeling.
Encoded Context Augmentation: Concatenating retrieved documents with the query becomes infeasible as the sequence length and sample size grow. Encoding the context and fusing it with the decoder (Fusion-in-Decoder) using cross-attention makes it possible to augment more samples without increasing computation costs significantly [257, 238, 255, 25].
Web Augmented: Locally stored memory, but external to LLM, has limited information. However, a large amount of information is available on the internet, which gets updated regularly. Rather than storing information locally, various methods retrieve query-related context through a web search and forward it to LLMs [258, 259, 157].
III-F2 Tool Augmented LLMs
While RAG relies on the retriever to provide context to the LLM to answer queries, tool augmented LLMs capitalize on the reasoning abilities of LLMs to iteratively plan by dividing tasks into sub-tasks, selecting necessary tools, and taking actions to complete the task [260, 261, 262, 27]. A generic pipeline of tool-augmented LLMs is shown in Figure 13, where different modules in Figure 13 are selected in a loop until the task completion.
Figure 13: A basic flow diagram of tool augmented LLMs. Given an input and a set of available tools, the model generates a plan to complete the task. The tool augmented LLMs utilize different modules iteratively, such as retriever, tool execution, read-write to memory, feedback, etc., depending on the task.
Zero-Shot Tool Augmentation: LLMs in-context learning and reasoning abilities enable them to interact with tools without training. Automatic reasoning and tool-use (ART) [262] builds a task library with demonstrations of reasoning steps and calling external tools. It retrieves similar task examples and provides the context to the LLM for inference. Aside from this, [263] shows tool documentation is enough to teach LLMs to use tools without demonstrations. RestGPT [264] integrates LLMs with RESTful APIs by decomposing tasks into planning and API selection steps. The API selector understands the API documentation to select a suitable API for the task and plan the execution. ToolkenGPT [265] uses tools as tokens by concatenating tool embeddings with other token embeddings. During inference, the LLM generates the tool tokens representing the tool call, stops text generation, and restarts using the tool execution output.
Training with Tool Augmentation: LLMs are trained to interact with diverse tools, enhancing planning abilities to overcome the limitations of zero-shot tool augmentation [266, 27, 267, 268]. Gorilla [266] instruction-tunes LLaMA with information retrieval from API documentation. It uses self-instruct [19] data generation pipeline with GPT-4 by providing in-context examples retrieved from API documentation. Tool augmented language model (TALM) [27] fine-tunes T5 [10] for tool use with a self-play approach, where it iteratively completes tool manipulation tasks and includes them back in the training set. ToolLLM [268] collects 16k APIs from RapidAPI. It samples APIs from the list to generate an instruction-tuning dataset using ChatGPT in single-tool and multi-tool scenarios. For high-quality datasets, ToolLLM suggested a depth-first search-based decision tree (DFSDT) method to generate ground-truths with diverse reasoning and planning.
[…]
Figure 14: Illustration of parameter-efficient fine-tuning paradigms, where $x$ is input and $h$ is hidden state, figure courtesy [38].
III-G Efficient LLMs
III-G1 Parameter Efficient Fine-Tuning
Fine-tuning LLMs with tens or hundreds of billions of parameters, such as GPT-3 (175B), BLOOM (176B), MT-NLG (540B), etc., is hardware-extensive and time-consuming. To avoid complete model fine-tuning, numerous parameter-efficient fine-tuning (PEFT) [40, 273, 41, 38, 39] are trying to reach full model fine-tuning performance at reduced costs. PEFT outperforms low-resource tasks, achieves comparable performance on medium-resource tasks, and underperforms on high-resource tasks compared to full fine-tuning [274]. An overview of different PEFT approaches is shown in Figure 14.
Adapter Tuning: Adds a few trainable parameters within the transformer block. The adapter layer is a sequence of feature downscaling, non-linearity, and upscaling [102]. Variants of adapter tuning inject adapter layers sequentially [102] and parallel [275], whereas the mixture of adapter (AdaMix) [276] employs multiple adapter modules in a single layer. AdaMix routes input instances randomly to one of the multiple downscale and upscale modules. The mixture of adapters is averaged out for inference to avoid additional latency. Low-Rank Adaptation (LoRA) [233] learns low-rank decomposed matrices to freeze original weights. Learned weights are fused with the original weights for inference, avoiding latency.
Prompt Tuning: Prompting is an effective way to adapt a pre-trained LLM for the downstream task. However, manual prompts bring uncertainty in the model’s prediction, where a change in a single word drops the performance [273]. Prompt tuning alleviates this problem by fine-tuning only 0.001%-3% additional parameters [277]. It concatenates trainable prompt parameters with the model embeddings [273, 40, 277]. Task-specific fixed discrete prompts are concatenated with input embeddings in [40]. As discrete prompts bring instability, prompts are encoded through a learnable mapping in P-Tuning [273], naming continuous prompts, which are appended with the discrete prompts. Only the prompt encoder is trainable in the model. In an extension of P-Tuning, continuous prompts are concatenated with each layer of the network in [277]. Progressive prompts [278] avoid catastrophic forgetting and transfer previously learned knowledge by sequentially adding trainable prompt embeddings to the previously frozen task embeddings.
Prefix Tuning: A set of trainable task-specific prefix vectors are appended to the frozen transformer layers in prefix tuning [41]. The prefix vectors are virtual tokens attended by the context tokens on the right. In addition, adaptive prefix tuning [279] applies a gating mechanism to control the information from the prefix and actual tokens.
Bias Tuning: Fine-tuning only bias terms in small to medium training data has been found effective in BitFit [280]. This method achieves full fine-tuning performance for tasks with less training data and comparable performance with larger training data.
III-G2 Quantization
LLMs require extensive computing and memory for inference. Deploying the GPT-3 175B model needs at least 5x80GB A100 GPUs and 350GB of memory to store in FP16 format [281]. Such demanding requirements for deploying LLMs make it harder for smaller organizations to utilize them. Model compression is an effective solution but comes at the cost of degrading performance, especially at large scales greater than 6B. These models exhibit very large magnitude outliers that do not exist in smaller models [282], making it challenging and requiring specialized methods for quantizing LLMs [281, 283].
Post-Training Quantization: Minimal or no training is required in this quantization type without compromising the model’s performance. LLM-8-bit [282] uses full-precision matrix multiplication for weights associated with outlier features and 8-bit for remaining. The lower precision multiplication outputs are converted to FP-16 and concatenated with others. The quantized models have homogenous word embeddings, degrading their performance. To fix this, token-level knowledge distillation is employed in [284] along with independent quantization scaling factors for each module due to varying weight distribution. Feature distributions are asymmetric and appear in different channels; outlier suppresion+ [285] shifts and scales per-channel activation distributions for effective quantization. SmoothQuant [281] quantizes activations and weights to INT8 by smoothing activations and migrating the quantization difficulty toward weights. It multiplies the inverse of the smoothing factor with weights, which introduces a few outliers in the weights but is easier to quantify than unsmoothed activations. OPTQ [283] uses the optimal brain compression (OBC) [286] algorithm to quantize the model layer-by-layer and update weights to compensate for quantization error. To improve speed and performance, OPTQ updates weights in arbitrary order, employs lazy updates, and uses better Cholesky kernels. Outlier-aware weight quantization (OWQ) [287] uses the OPTQ algorithm for quantization but assigns higher precision to vulnerable weights, causing outliers and lower precision for others.
Quantization-Aware Training: To compensate for the performance degradation, the quantized model is fine-tuned in quantization-aware training (QAT) [288, 289, 290]. Alphatuning quantizes the model using binary coding quantization (BCQ) [291] and fine-tunes only quantization scaling factors. This approach improves performance over parameter-efficient fine-tuning of a pre-trained model and then quantizing and fine-tuning. Similarly, parameter-efficient and quantization-aware adaptation (PEQA) [292] reduces the precision of fully-connected layers and fine-tunes only quantization scaling parameters. LLM-QAT [290] generates training data from the pre-trained network and trains a quantized student model with knowledge distillation. QLoRA [289] fine-tunes 4-bit quantized pre-trained LLM with LoRA [233] using 4-bit normal float, which shows better performance over 4-bit integer and float.
III-G3 Pruning
Pruning is an alternative approach to quantization to compress model size, thereby reducing LLMs deployment costs significantly. Compared to task-agnostic pruning, task-specific pruning is easily achievable with good performance, where a model is fine-tuned on the downstream task and pruned for faster inference. It is possible to prune LLMs for individual tasks, but the cost of pruning and deploying task-specific models is high. To overcome this, many structured and unstructured pruning methods for LLMs have been proposed to maintain reasonable performance while shrinking the model size [293, 294, 295].
Unstructured Pruning: This kind of pruning removes less important weights without maintaining any structure. Existing LLM pruning methods take advantage of the unique characteristics of LLMs, uncommon for smaller models, where a small subset of hidden states are activated with large magnitude [282]. Pruning by weights and activations (Wanda) [293] prunes weights in every row based on importance, calculated by multiplying the weights with the norm of input. The pruned model does not require fine-tuning, saving large models’ computational costs. Outlier weighed layerwise sparsity (OWL) [296] extends Wanda with non-uniform layer pruning. It shows that the number of outliers varies for different layers; therefore, the model should have variable pruning ratios for better performance for every layer. Contrastive pruning (CAP) [297] iteratively prunes the model by training the sparse model using contrastive loss between pre-trained, fine-tuning, and snapshots of previous sparse models to learn task-specific and task-agnostic knowledge.
Structured Pruning: Here, the parameters are removed in groups, rows, columns, or matrices, which speeds up the inference because of effective NVIDIA tensor core utilization [293]. LLM-Pruner [294] employs a 3-stage structured pruning strategy, identifying the groups of hidden states causing each other to activate during forward-pass, keeping important groups and removing less important ones, and fine-tuning the pruned model with LoRA. Sparsity-induced mask learning (SIMPLE) [298] prunes the network using learnable masks. Similarly, another method prunes LLMs by learning masks and removing unimportant rank-1 components of the factorized weight matrix [295].
TABLE III: Summary of pre-trained LLMs (>10B). Only the LLMs discussed individually in the previous sections are summarized. “Data/Tokens” is the model’s pre-training data, which is either the number of tokens or data size. “Data Cleaning” indicates whether the data cleaning is performed or not. This includes heuristics (Heur), deduplication (Dedup), quality filtering (QF), and privacy filtering (PF), “Cost” is the calculated training cost obtained by multiplying the GPUs/TPUs hourly rate with the number of GPUs and the training time. The actual cost may vary due to many reasons such as using in-house GPUs or getting a discounted rate, re-training, number of employees working on the problem, etc. “Training Parallelism” indicates distributed training using data parallelism (D), tensor parallelism (T), pipeline parallelism (P), model parallelism (M), optimizer parallelism (OP), and rematerialization (R), where for “Library” column, “DS” is a short form for Deep Speed. In column “Commercial Use”, we assumed a model is for non-commercial purposes if its license is unavailable.
Models Publication Venue License Type Model Creators Purpose No. of Params   Commercial Use Steps Trained Data/ Tokens Data Cleaning No. of Processing Units Processing Unit Type Training Time Calculated Train. Cost Training Parallelism Library
T5 [10] JMLR'20 Apache-2.0 Google General 11B $\checkmark$ 1M 1T Heur+Dedup 1024 TPU v3 - - D+M Mesh TensorFlow
GPT-3 [6] NeurIPS'20 - OpenAI General 175B $\times$ - 300B Dedup+QF - V100 - - M -
mT5 [11] NAACL'21 Apache-2.0 Google General 13B $\checkmark$ 1M 1T - - - - - - -
PanGu-$\alpha$ [104] arXiv'21 Apache-2.0 Huawei General 200B $\checkmark$ 260k 1.1TB Heur+Dedup 2048 Ascend 910 - - D+OP+P+O+R MindSpore
CPM-2 [12] AI Open'21 MIT Tsinghua General 198B $\checkmark$ 1M 2.6TB Dedup - - - - D+M JAXFormer
Codex [131] arXiv'21 - OpenAI Coding 12B $\times$ - 100B Heur - - - - - -
ERNIE 3.0 [107] arXiv'21 - Baidu General 10B $\times$ $*{}^{*}$ 375B Heur+Dedup 384 V100 - - $*{}^{*}$ PaddlePaddle
Jurassic-1 [109] White-Paper'21 Apache-2.0 AI21 General 178B $\checkmark$ - 300B - 800 GPU - - D+M+P Megatron+DS
HyperCLOVA [111] EMNLP'21 - Naver General 82B $\times$ - 300B Clf+Dedup+PF 1024 A100 321h 1.32 Mil M Megatron
Yuan 1.0 [112] arXiv'21 Apache-2.0 - General 245B $\checkmark$ $*{}^{*}$ 180B Heur+Clf+Dedup 2128 GPU - - D+T+P -
Gopher [113] arXiv'21 - Google General 280B $\times$ - 300B QF+Dedup 4096 TPU v3 920h 13.19 Mil D+M JAX+Haiku
ERNIE 3.0 Titan [35] arXiv'21 - Baidu General 260B $\times$ - 300B Heur+Dedup - Ascend 910 - - D+M+P+D* PaddlePaddle
GPT-NeoX-20B [299] BigScience'22 Apache-2.0 EleutherAI General 20B $\checkmark$ 150k 825GB None 96 40G A100 - - M Megatron+DS+PyTorch
OPT [14] arXiv'22 MIT Meta General 175B $\checkmark$ 150k 180B Dedup 992 80G A100 - - D+T Megatron
BLOOM [13] arXiv'22 RAIL-1.0 BigScience General 176B $\checkmark$ - 366B Dedup+PR 384 80G A100 2520h 3.87 Mil D+T+P Megatron+DS
Galactica [138] arXiv'22 Apache-2.0 Meta Science 120B $\times$ 225k 106B Dedup 128 80GB A100 - - - Metaseq
GLaM [118] ICML'22 - Google General 1.2T $\times$ $*{}^{*}$ 600B Clf 1024 TPU v4 - - M GSPMD
LaMDA [140] arXiv'22 - Google Dialog 137B $\times$ 3M 2.81T Filtered 1024 TPU v3 1384h 4.96 Mil D+M Lingvo
MT-NLG [114] arXiv'22 Apache-v2.0 MS.+Nvidia General 530B $\times$ - 270B - 4480 80G A100 - - D+T+P Megatron+DS
AlphaCode [132] Science'22 Apache-v2.0 Google Coding 41B $\checkmark$ 205k 967B Heur+Dedup - TPU v4 - - M JAX+Haiku
Chinchilla [121] arXiv'22 - Google General 70B $\times$ - 1.4T QF+Dedup - TPUv4 - - - JAX+Haiku
PaLM [15] arXiv'22 - Google General 540B $\times$ 255k 780B Heur 6144 TPU v4 - - D+M JAX+T5X
AlexaTM [122] arXiv'22 Apache v2.0 Amazon General 20B $\times$ 500k 1.1T Filtered 128 A100 2880h 1.47 Mil M DS
U-PaLM [124] arXiv'22 - Google General 540B $\times$ 20k - - 512 TPU v4 120h 0.25 Mil - -
UL2 [89] ICLR'23 Apache-2.0 Google General 20B $\checkmark$ 2M 1T - 512 TPU v4 - - M JAX+T5X
GLM [33] ICLR'23 Apache-2.0 Multiple General 130B $\times$ - 400B - 768 40G A100 1440h 3.37 Mil M -
CodeGen [130] ICLR'23 Apache-2.0 Salesforce Coding 16B $\checkmark$ 650k 577B Heur+Dedup - TPU v4 - - D+M JAXFormer
LLaMA [126] arXiv'23 - Meta General 65B $\times$ 350k 1.4T Clf+Heur+Dedup 2048 80G A100 504h 4.12 Mil D+M xFormers
[…]
BloombergGPT [141] arXiv23 - Bloomberg Finance 50B $\times$ 139k 569B Dedup 512 40G A100 1272h 1.97 Mil M PyTorch
Xuan Yuan 2.0 [143] arXiv23 RAIL-1.0 Du Xiaoman Finance 176B $\checkmark$ - 366B Filtered 80GB A100 - - P DS
CodeT5+ [34] arXiv'23 BSD-3 Salesforce Coding 16B $\checkmark$ 110k 51.5B Dedup 16 40G A100 - - - DS
StarCoder [137] arXiv'23 OpenRAIL-M BigCode Coding 15.5B $\checkmark$ 250k 1T Dedup+QF+PF 512 80G A100 624h 1.28 Mil D+T+P Megatron-LM
LLaMA-2 [21] arXiv'23 LLaMA-2.0 Meta General 70B $\checkmark$ 500k 2T Minimal Filtering - 80G A100 1.7Mh - - -
PaLM-2 [123] arXiv'23 - Google General - $\times$ - - Ddedup+PF+QF - - - - - -
TABLE IV: Summary of instruction tuned LLMs (>10B). All abbreviations are the same as Table III. Entries in “Data/Tokens” starting with “S-” represents the number of training samples.
Models Publication Venue License Type Model Creators Purpose No. of Params Commercial Use Pre-trained Models Steps Trained Data/ Tokens No. of Processing Units Processing Unit Type Train. Time Calculated Train. Cost Train. Parallelism Library
WebGPT [157] arXiv'21 - OpenAI General 175B $\times$ GPT-3 - - - - - - - -
T0 [17] ICLR'22 Apache-2.0 BigScience General 11B $\checkmark$ T5 - 250B 512 TPU v3 270h 0.48 Mil - -
Tk-Instruct [18] EMNLP'22 MIT AI2+ General 11B $\checkmark$ T5 1000 - 256 TPU v3 4h 0.0036 Mil - Google T5
OPT-IML [93] arXiv'22 - Meta General 175B $\times$ OPT 8k 2B 128 40G A100 - - D+T Megatron
Flan-U-PaLM [16] ICLR'22 Apache-2.0 Google General 540B $\checkmark$ U-PaLM 30k - 512 TPU v4 - - - JAX+T5X
mT0 [145] ACL'23 Apache-2.0  HuggingFace+ General 13B $\checkmark$ mT5 - - - - - - - -
Sparrow [158] arXiv'22 - Google Dialog 70B $\times$ Chinchilla - - 64 TPU v3 - - M -
WizardCoder [155] arXiv'23 Apache-2.0 HK Bapt. Coding 15B $\times$ StarCoder 200 S-78k - - - - - -
Alpaca [149] Github'23 Apache-2.0 Stanford General 13B $\checkmark$ LLaMA 3-Epoch S-52k 8 80G A100 3h 600 FSDP PyTorch
Vicuna [150] Github'23 Apache-2.0 LMSYS General 13B $\checkmark$ LLaMA 3-Epoch S-125k - - - - FSDP PyTorch
LIMA [176] arXiv'23 - Meta+ General 65B - LLaMA 15-Epoch S-1000 - - - - - -
Koala [300] Github'23 Apache-2.0 UC-Berkley General 13B $\times$ LLaMA 2-Epoch S-472k 8 A100 6h 100 - JAX/FLAX
IV Findings & Insights
Training a billion-scale model is difficult as compared to a smaller model. LLMs are prone to various instabilities during training, such as hardware failure and instability. Other than this, LLMs exhibit different behaviors such as emergent abilities, improved zero-shot, few-shot, and reasoning abilities. Researchers report these essential details in their papers for results reproduction and field progress. We identify critical information in Table I and II such as architecture, training strategies, and pipelines that improve LLMs’ performance or other abilities acquired because of changes mentioned in section III.
V Model Configurations
We provide different statistics of pre-trained and instruction-tuned models in this section. This includes information such as publication venue, license type, model creators, steps trained, parallelism, etc in Table III and Table IV. Architecture details of pre-trained LLMs are available in Table V. Providing these details for instruction-tuned models is unnecessary because it fine-tunes pre-trained models for instruction datasets. Hence, architectural details are the same as the baselines. Moreover, optimization settings for various LLMs are available in Table VI and Table VII. We do not include details on precision, warmup, and weight decay in Table VII. Neither of these details are important as others to mention for instruction-tuned models nor provided by the papers.
TABLE V: Architecture details of LLMs. Here, “PE” is the positional embedding, “nL” is the number of layers, “nH” is the number of attention heads, “HS” is the size of hidden states.
Models Type
Training
Objective
Attention Vocab Tokenizer Norm PE Activation Bias nL nH HS
T5 (11B) Enc-Dec Span Corruption Standard 32k SentencePiece Pre-RMS Relative ReLU $\times$ 24 128 1024
GPT3 (175B) Causal-Dec Next Token Dense+Sparse - - Layer Learned GeLU 96 96 12288
mT5 (13B) Enc-Dec Span Corruption Standard 250k SentencePiece Pre-RMS Relative ReLU - - - -
PanGu-$\alpha$ (200B) Causal-Dec Next Token Standard 40k BPE Layer - - - 64 128 16384
CPM-2 (198B) Enc-Dec Span Corruption Standard 250k SentencePiece Pre-RMS Relative ReLU - 24 64 -
Codex (12B) Causal-Dec Next Token Standard - BPE+ Pre-Layer Learned GeLU - 96 96 12288
ERNIE 3.0 (10B) Causal-Dec Next Token Standard - WordPiece Post-Layer Relative GeLU - 48 64 4096
Jurassic-1 (178B) Causal-Dec Next Token Standard 256k $*{}^{*}$ Pre-Layer Learned GeLU $\checkmark$ 76 96 13824
HyperCLOVA (82B) Causal-Dec Next Token Dense+Sparse - BPE* Pre-Layer Learned GeLU - 64 80 10240
Yuan 1.0 (245B) Causal-Dec Next Token Standard - - - - - - 76 - 16384
Gopher (280B) Causal-Dec Next Token Standard 32k SentencePiece Pre-RMS Relative GeLU $\checkmark$ 80 128 16384
ERNIE 3.0 Titan (260B) Causal-Dec Next Token Standard - WordPiece Post-Layer Relative GeLU - 48 192 12288
GPT-NeoX-20B Causal-Dec Next Token Parallel 50k BPE Layer Rotary GeLU $\checkmark$ 44 64 -
OPT (175B) Causal-Dec Next Token Standard - BPE - - ReLU $\checkmark$ 96 96 -
BLOOM (176B) Causal-Dec Next Token Standard 250k BPE Layer ALiBi GeLU $\checkmark$ 70 112 14336
Galactica (120B) Causal-Dec Next Token Standard 50k BPE+custom Layer Learned GeLU $\times$ 96 80 10240
GLaM (1.2T) MoE-Dec Next Token Standard 256k SentencePiece Layer Relative GeLU $\checkmark$ 64 128 32768
LaMDA (137B) Causal-Dec Next Token Standard 32k BPE Layer Relative GeGLU - 64 128 8192
MT-NLG (530B) Causal-Dec Next Token Standard 50k BPE Pre-Layer Learned GeLU $\checkmark$ 105 128 20480
AlphaCode (41B) Enc-Dec Next Token Multi-query 8k SentencePiece - - - - 64 128 6144
Chinchilla (70B) Causal-Dec Next Token Standard 32k SentencePiece-NFKC Pre-RMS Relative GeLU $\checkmark$ 80 64 8192
PaLM (540B) Causal-Dec Next Token Parallel+Multi-query 256k SentencePiece Layer RoPE SwiGLU $\times$ 118 48 18432
AlexaTM (20B) Enc-Dec Denoising Standard 150k SentencePiece Pre-Layer Learned GeLU $\checkmark$ 78 32 4096
Sparrow (70B) Causal-Dec Pref.&Rule RM - 32k SentencePiece-NFKC Pre-RMS Relative GeLU $\checkmark$ $*{}^{*}$ 64 8192
U-PaLM (540B) Non-Causal-Dec MoD Parallel+Multi-query 256k SentencePiece Layer RoPE SwiGLU $\times$ 118 48 18432
UL2 (20B) Enc-Dec MoD Standard 32k SentencePiece - - - - 64 16 4096
GLM (130B) Non-Causal-Dec AR Blank Infilling Standard 130k SentencePiece Deep RoPE GeGLU 70 96 12288
CodeGen (16B) Causal-Dec Next Token Parallel - BPE Layer RoPE - - 34 24 -
LLaMA (65B) Causal-Dec Next Token Standard 32k BPE Pre-RMS RoPE SwiGLU - 80 64 8192
PanGu-$\Sigma$ (1085B) Causal-Dec Next Token Standard - BPE Fused Layer - FastGeLU - 40 40 5120
BloombergGPT (50B) Causal-Dec Next Token Standard 131k Unigram Layer ALiBi GeLU $\checkmark$ 70 40 7680
Xuan Yuan 2.0 (176B) Causal-Dec Next Token Self 250k BPE Layer ALiBi GeLU $\checkmark$ 70 112 14336
CodeT5+ (16B) Enc-Dec SC+NT+Cont.+Match Standard - Code-Specific - - - - - - -
StarCoder (15.5B) Causal-Dec FIM Multi-query 49k BPE - Learned - - 40 48 6144
LLaMA (70B) Causal-Dec Next Token Grouped-query 32k BPE Pre-RMS RoPE SwiGLUE - - - -
PaLM-2 - MoD Parallel - - - - - - - - -
TABLE VI: Summary of optimization settings used for pre-trained LLMs. The values for weight decay, gradient clipping, and dropout are 0.1, 1.0, and 0.1, respectively, for most of the LLMs.
Sequence LR Optimizers Precision Weight Grad
Models Batch Size Length LR Warmup Decay AdaFactor Adam AdamW FP16 BF16 Mixed Decay Clip Dropout
[…]
$\checkmark$
- - $\checkmark$
GPT3 (175B) 32K - 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
mT5 (13B) 1024 1024 0.01 - inverse square root
$\checkmark$
- - $\checkmark$
PanGu-$\alpha$ (200B) - 1024 2e-5 - -
$\checkmark$
CPM-2 (198B) 1024 1024 0.001 - -
$\checkmark$
- - $\checkmark$
Codex (12B) - - 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ - -
ERNIE 3.0 (12B) 6144 512 1e-4 $\checkmark$ linear
$\checkmark$
$\checkmark$ - -
Jurassic-1 (178B) 3.2M 2048 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
HyperCLOVA (82B) 1024 - 6e-5 - cosine
$\checkmark$
$\checkmark$ - -
Yuan 1.0 (245B) $<$10M 2048 1.6e-4 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$ - -
Gopher (280B) 3M 2048 4e-5 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$
- $\checkmark$ -
ERNIE 3.0 Titan (260B) - 512 1e-4 $\checkmark$ linear
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
GPT-NeoX-20B 1538 2048 0.97e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\times$
OPT (175B) 2M 2048 1.2e-4 - linear
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\checkmark$
BLOOM (176B) 2048 2048 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\times$
Galactica (120B) 2M 2048 7e-6 $\checkmark$ linear decay to 10%
$\checkmark$
$\checkmark$ $\checkmark$ $\checkmark$
GLaM (1.2T) 1M 1024 0.01 - inverse square root
$\checkmark$
FP32 + $\checkmark$ - $\checkmark$ $\times$
LaMDA (137B) 256K - - - -
MT-NLG (530B) 1920 2048 5e-5 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
AlphaCode (41B) 2048 1536+768 1e-4 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
Chinchilla (70B) 1.5M 2048 1e-4 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$
PaLM (540B) 2048 2048 0.01 - inverse square root
$\checkmark$
$\checkmark$ $\checkmark$ $\times$
AlexaTM (20B) 2M 1024 1e-4 - linear decay to 5%
$\checkmark$
$\checkmark$
$\checkmark$ - $\checkmark$
U-PaLM (540B) 32 2048 1e-4 - cosine
$\checkmark$
UL2 (20B) 1024 1024 - - inverse square root
$\times$ - -
GLM (130B) 4224 2048 8e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\checkmark$
CodeGen (16B) 2M 2048 5e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$ $\checkmark$ -
LLaMA (65B) 4M Tokens 2048 1.5e-4 $\checkmark$ cosine decay to 10%
$\checkmark$
$\checkmark$ $\checkmark$ -
PanGu-$\Sigma$ (1.085T) 512 1024 2e-5 $\checkmark$ -
$\checkmark$
$\checkmark$
BloombergGPT (50B) 2048 2048 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ $\times$
Xuan Yuan 2.0 (176B) 2048 2048 6e-5 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
CodeT5+ (16B) 2048 1024 2e-4 - linear
$\checkmark$
$\checkmark$
$\checkmark$ - -
StarCoder (15.5B) 512 8k 3e-4 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ - -
LLaMA-2 (70B) 4M Tokens 4k 1.5e-4 $\checkmark$ cosine
$\checkmark$
$\checkmark$
$\checkmark$ $\checkmark$ -
TABLE VII: Summary of optimization settings used for instruction-tuned LLMs. Values for gradient clipping and dropout are the same as the pre-trained models, while no model uses weight decay for instruction tuning.
Sequence Optimizers Grad
Models Batch Size Length LR Warmup LR_Decay AdaFactor Adam AdamW Clip Dropout
WebGPT (175B) BC:512, RM:32 - 6e-5 - - $\checkmark$ - -
T0 (11B) 1024 1280 1e-3 - - $\checkmark$ - $\checkmark$
Tk-Instruct (11B) 1024 - 1e-5 - constant - - - - -
OPT-IML (175B) 128 2048 5e-5 $\times$ linear $\checkmark$ $\checkmark$ $\checkmark$
Flan-U-PaLM (540B) 32 - 1e-3 - constant $\checkmark$ - $\checkmark$
Sparrow (70B) RM: 8+16, RL:16 - 2e-6 $\checkmark$ cosine decay to 10% $\checkmark$ $\checkmark$ $\times$
WizardCoder (15B) 512 2048 2e-5 $\checkmark$ cosine - - - - -
Alpaca (13B) 128 512 1e-5 $\checkmark$ cosine - - $\checkmark$ $\checkmark$ $\times$
Vicuna (13B) 128 -2048 2e-5 $\checkmark$ cosine $\checkmark$ - $\times$
LIMA (65B) 32 2048 1e-5 $\times$ linear $\checkmark$ - $\checkmark$
VI Datasets and Evaluation
Generating training and evaluation datasets is expensive because of the large-scale data demand of LLMs. Hence, datasets for training and benchmarking these models are topics of key importance. In Fig. 15, we show the distribution of the existing datasets for various NLP tasks. We restrict our distribution to only the most important tasks in the literature by including tasks with at least 20 datasets. LLMs can directly benefit from these datasets for training and evaluation. A summary of the training and evaluation datasets commonly used by LLMs is provided next.
VI-A Training Datasets
The performance of LLMs largely depends on the training data’s quality, size, and diversity. Preparing training datasets of high quality at a large scale is laborious. Researchers have suggested various pre-training and fine-tuning datasets to enhance LLMs capabilities. We summarize these efforts in Table VIII. While numerous training datasets are available in the literature, we cover the most widely used ones in our summary.
VI-B Evaluation Datasets and Tasks
The evaluation of LLMs is important in gauging their proficiency and limitations. This process measures the model’s ability to comprehend, generate, and interact with human language across a spectrum of tasks. Evaluating a language model (LM) is divided into two broader categories: 1) natural language understanding (NLU) and 2) natural language generation (NLG). It is emphasized that tasks in NLU and NLG are softly categorized and are often used interchangeably in the literature.
Natural Language Understanding: This task measures the language understanding capacity of LMs. It encompasses multiple tasks, including sentiment analysis, text classification, natural language inference (NLI), question answering (QA), commonsense reasoning (CR), mathematical reasoning (MR), reading comprehension (RC), etc.
Natural Language Generation: This task assesses the language generation capabilities of LLMs by understanding the provided input context. It includes tasks such as summarization, sentence completion, machine translation (MT), dialogue generation, etc.
Numerous datasets are proposed for each task, evaluating LLMs against different characteristics. To provide an overview of evaluation datasets, we briefly discuss a few famous datasets within each category and offer a comprehensive list of datasets in Table IX. Moreover, we show a detailed overview of the training datasets and evaluation tasks and benchmarks used by various pre-trained LLMs in Table X and fine-tuned LLMs in Table XI. We also compare the top-performing LLMs in various NLP tasks in Table XII.
TABLE VIII: Details of various well-known pre-training and fine-tuning datasets. Here, alignment means aligning with human preferences.
Dataset Type Size/Samples Tasks Source Creation
Comments
C4 [10] Pretrain 806GB - Common Crawl Automated
A clean, multilingual dataset with billions of tokens
mC4 [11] Pretrain 38.49TB - Common Crawl Automated
A multilingual extension of the C4 dataset, mC4 identifies over 100 languages using cld3 from 71 monthly web scrapes of Common Crawl.
PILE [301] Pretrain 825GB -
Common Crawl, PubMed Central,
OpenWebText2, ArXiv, GitHub,
Books3, and others
Automated
A massive dataset comprised of 22 constituent sub-datasets
ROOTs [302] Pretrain 1.61TB - 498 Hugging Face datasets Automated
46 natural and 13 programming languages
MassiveText [113] Pretrain 10.5TB -
MassiveWeb, Books, News,
Wikipedia, Github, C4
Automated
99% of the data is in English
Wikipedia [303] Pretrain - - Wikipedia Automated
Dump of wikipedia
RedPajama [304] Pretrain 5TB -
CommonCrawl, C4, Wikipedia,
Github, Books, StackExchange
Automated
Open-source replica of LLaMA dataset
PushShift.io Reddit Pretrain 21.1GB - Reddit Automated
Submissions and comments on Reddit from 2005 to 2019
BigPython [130] Pretrain 5.5TB Coding GitHub Automated
Pool of Prompt (P3) [17] Instructions 12M 62 PromptSource Manual
A Subset of PromptSource, created from 177 datasets including summarization, QA, classification, etc.
xP3 [145] Instructions 81M 71 P3+Multilingual datasets Manual
Extending P3 to total 46 languages
Super-NaturalInstructions (SNI) [18] Instructions 12.4M 1616 Multiple datasets Manual
Extending P3 with additional multi-lingual datasets, total 46 languages
Flan [16] Instructions 15M 1836 Muffin+T0-SF+NIV2 Manual
Total 60 languages
OPT-IML [93] Instructions 18.1M 1667 - Manual
Self-Instruct [19] Instructions 82k 175 - Automated
Generated 52k instructions with 82k samples from 175 seed tasks using GPT-3
Alpaca [149] Instructions 52k - - Automated
Employed self-instruct method to generate data from text-davinci-003
Vicuna [150] Instructions 125k - ShareGPT Automated
Conversations shared by users on ShareGPT using public APIs
LLaMA-GPT-4 [151] Instructions 52k - Alpaca Automated
Recreated Alpaca dataset with GPT-4 in English and Chinese
Unnatural Instructions [305] Instructions 68k - 15-Seeds (SNI) Automated
LIMA [176] Instructions 1k - Multiple datasets Manual
Carefully created samples to test performance with fine-tuning on less data
Anthropic-HH-RLHF [306] Alignment 142k - - Manual
Anthropic-HH-RLHF-2 [169] Alignment 39k - - Manual
Figure 15: A distribution of datasets proposed for different NLP tasks. We include only the tasks for which at least 20 datasets have already been proposed.
TABLE IX: Categorized evaluation datasets used in evaluating LLMs.
Type Datasets/Benchmarks
Multi-Task
MMLU [307], SuperGLUE [2], BIG-bench [308], GLUE [309], BBH [308], CUGE [310], ZeroCLUE [311], FewCLUE [312], Blended Skill Talk [313], HELM [314], KLUE-STS [315]
Language Understanding
CoQA [316], WiC [317], Wikitext103 [318], PG19 [319], LCQMC [320], QQP [321], WinoGender [322], CB [323], FinRE [324], SanWen [325], AFQMC [311], BQ Corpus [326], CNSS [327], CKBQA 13 [328], CLUENER [311], Weibo [329], AQuA [330], OntoNotes [331], HeadQA [332], Twitter Dataset [333]
Story Cloze and
Sentence Completion
StoryCloze [334], LAMBADA [335], LCSTS [336], AdGen [337], E2E [338], CHID [339], CHID-FC [312]
Physical Knowledge and
World Understanding
PIQA [340], TriviaQA [341], ARC [342], ARC-Easy [342], ARC-Challenge [342], PROST [343], OpenBookQA [344], WebNLG [345], DogWhistle Insider & Outsider [346]
Contextual Language
Understanding
RACE [347], RACE-Middle [347], RACE-High [347], QuAC [348], StrategyQA [349], Quiz Bowl [350],
cMedQA [351], cMedQA2 [352], MATINF-QA [353]
Commonsense Reasoning
WinoGrande [354], HellaSwag [355], COPA [356], WSC [357], CSQA [358], SIQA [359], C3 [360], CLUEWSC2020 [311], CLUEWSC [311], CLUEWSC-FC [312], ReCoRD [361]
Reading Comprehension
SQuAD [362], BoolQ [363], SQUADv2 [364], DROP [365], RTE [366], WebQA [367], CMRC2017 [368], CMRC2018 [369], CMRC2019 [370], COTE-BD [371], COTE-DP [371], COTE-MFW [371], MultiRC [372], Natural Questions [373], CNSE [327], DRCD [374], DuReader [375], Dureaderrobust [376], DuReader-QG [375], SciQ [377], Sogou-log [378], Dureaderrobust-QG [376], QA4MRE [379], KorQuAD 1.0 [380], CAIL2018-Task1 & Task2 [381]
Mathematical Reasoning
MATH [382], Math23k [383], GSM8K [384], MathQA [385], MGSM [386], MultiArith [387], ASDiv [388], MAWPS [389], SVAMP [390]
Problem Solving
HumanEval [391], DS-1000 [392], MBPP [393], APPS [382], CodeContests [132]
Natural Language Inference
& Logical Reasoning
ANLI [394], MNLI-m [395], MNLI-mm [395],QNLI [362], WNLI [357], OCNLI [311], CMNLI [311], ANLI R1 [394], ANLI R2 [394], ANLI R3 [394], HANS [396], OCNLI-FC [312], LogiQA [397], StrategyQA [349]
Cross-Lingual Understanding
MLQA [398], XNLI [399], PAWS-X [400], XSum [401], XCOPA [402], XWinograd [403], TyDiQA-GoldP [404], MLSum [405]
Truthfulness and Fact Checking
TruthfulQA [406], MultiFC [407], Fact Checking on Fever [408]
Biases and Ethics in AI
ETHOS [409], StereoSet [410], BBQ [411], Winobias [412], CrowS-Pairs [413]
Toxicity
RealToxicityPrompts [414], CivilComments toxicity classification [415]
Language Translation
WMT [416], WMT20 [417], WMT20-enzh [417], EPRSTMT [312], CCPM [418]
Scientific Knowledge
AminoProbe [138], BioLAMA [138], Chemical Reactions [138], Galaxy Clusters [138], Mineral Groups [138]
Dialogue
Wizard of Wikipedia [419], Empathetic Dialogues [420], DPC-generated [121] dialogues, ConvAI2 [421], KdConv [422]
Topic Classification
TNEWS-FC [312], YNAT [315], KLUE-TC [315], CSL [311], CSL-FC [312], IFLYTEK [423]
TABLE X: An illustration of training datasets and evaluation tasks employed by pre-trained LLMs. Here, “QA” is question-answering, “Clf” is classification, “NLI” is natural language inference, “MT” is machine translation, “RC” is reading comprehension, “CR” is commonsense reasoning, “MR” is mathematical reasoning, “Mem.” is memorization.
Benchmark
Models
Training Dataset
BIG-
bench
MMLU
Super
GLUE
QA Clf NLI MT
Cloze/
Completion
RC CR MR Coding
Truthful/
Bias/
Toxicity/
Mem.
T5
C4 [10]
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GPT-3
Common Crawl, WebText, Books Corpora, Wikipedia
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
mT5
mC4 [11]
$\checkmark$ $\checkmark$ $\checkmark$
PanGu-$\alpha$
1.1TB Chinese Text Corpus
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CPM-2
WuDaoCorpus [105]
$\checkmark$ $\checkmark$
Codex
54 million public repositories from Github
$\checkmark$
ERNIE-3.0
Chinese text corpora, Baidu Search, Web text, QA-long, QA-short, Poetry and Couplet Domain-specific data from medical, law, and financial area Baidu knowledge graph with more than 50 million facts
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Jurassic-1
Wikipedia, OWT, Books, C4, Pile [301], arXiv, GitHub
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
HyperCLOVA
Korean blogs, Community sites, News, KiN Korean Wikipedia, Wikipedia (English and Japanese), Modu-Corpus: Messenger, News, Spoken and written language corpus, Web corpus
$\checkmark$
Yuan 1.0
Common Crawl, SogouT, Sogou News, Baidu Baike, Wikipedia, Books
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Gopher
subsets of MassiveWeb Books, C4, News, GitHub and Wikipedia samples from MassiveText
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ERNIE-3.0 TITAN
Same as ERNIE 3.0 and ERNIE 3.0 adversarial dataset, ERNIE 3.0 controllable dataset
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GPT-NeoX-20B
Pile [301]
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OPT
RoBERTa [424], Pile [301], PushShift.io Reddit [425]
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BLOOM
ROOTs [13]
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Galactica
arXiv, PMC, Semantic Scholar, Wikipedia, StackExchange, LibreText, Open Textbooks, RefSeq Genome, OEIS, LIPID MAPS, NASAExoplanet, Common Crawl, ScientificCC, AcademicCC, GitHub repositories Khan Problems, GSM8K, OneSmallStep
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GLaM
Filtered Webpages, Social media conversations Wikipedia, Forums, Books, News
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LaMDA
Infiniset : Public documents, Dialogs, Utterances
$\checkmark$
MT-NLG
Two snapshots of Common Crawl and Books3, OpenWebText2, Stack Exchange, PubMed Abstracts, Wikipedia, PG-19 [242], BookCorpus2, NIH ExPorter, Pile, CC-Stories, RealNews
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AlphaCode
Selected GitHub repositories, CodeContests: Codeforces, Description2Code, CodeNet
$\checkmark$
Chinchilla
MassiveWeb, MassiveText Books, C4, News, GitHub, Wikipedia
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PaLM
webpages, books, Wikipedia, news, articles, source code, social media conversations
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AlexaTM
Wikipedia, mC4
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
U-PaLM
Same as PaLM
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
UL2
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GLM-130B
$\checkmark$ $\checkmark$ $\checkmark$
CodeGen
Pile, BigQuery, BigPython
$\checkmark$
LLaMA
CommonCrawl, C4, Github, Wikipedia, Books, arXiv, StackExchange
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PanGu-$\Sigma$
WuDaoCorpora, CLUE, Pile, C4, Python code
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BloombergGPT
inPile, Pile, C4, Wikipedia
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CodeT5+
CodeSearchNet, Github Code
$\checkmark$ $\checkmark$
StarCoder
The Stack v1.2
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LLaMA-2
$\checkmark$
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PaLM-2
Web documents, Code, Books, Maths, Conversation
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TABLE XI: An illustration of training datasets and evaluation benchmarks used in fine-tuned LLMs. “SNI” is a short of Super-NaturalInsturctions.
Models
Training Dataset
BIG-
bench
MMLU BBH RAFT FLAN SNI PromptSource TyDiQA HumanEval MBPP
Truthful/
Bias/
Toxicity
T0
Pool of Prompts
$\checkmark$
WebGPT
ELI5 [426], ELI5 fact-check [157], TriviaQA [341], ARC-Challenge [342], ARC-Easy [342], Hand-written data, Demonstrations of humans, Comparisons between model-generated answers
$\checkmark$
Tk-INSTRUCT
SNI [18]
$\checkmark$
mT0
xP3 [145]
OPT-IML
PromptSource [17], FLAN [16], SNI [427], UnifiedSKG [428], CrossFit [429], ExMix [430], T5 [10], Reasoning
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Flan
Muffin, T0-SF, NIv2, CoT
$\checkmark$ $\checkmark$ $\checkmark$
WizardCoder
Code Alpaca
$\checkmark$ $\checkmark$
VI-B1 Multi-task
MMLU [307]
A benchmark that measures the knowledge acquired by models during pretraining and evaluates models in zero-shot and few-shot settings across 57 subjects, testing both world knowledge and problem-solving ability.
SuperGLUE [2]
A more challenging and diverse successor to the GLUE [309] benchmark, SuperGLUE includes a variety of language understanding tasks, such as question answering, natural language inference, and coreference resolution. It is designed to provide a rigorous test of language understanding and requires significant progress in areas like sample-efficient, transfer, multitasking, and unsupervised or self-supervised learning.
BIG-bench [308]
The BIG-bench (Behavior of Intelligent Generative Models Benchmark) is a large-scale benchmark designed to test the abilities of LLMs across a wide range of tasks, including reasoning, creativity, ethics, and understanding of specific domains.
GLUE [309]
The General Language Understanding Evaluation (GLUE) benchmark is a collection of resources for training, evaluating, and analyzing natural language understanding systems. It includes a variety of tasks that test a wide range of linguistic phenomena, making it a comprehensive tool for evaluating language understanding in AI.
VI-B2 Language Understanding
WinoGrande [354]
A large-scale dataset inspired by the original Winograd [357] Schema Challenge tests models on their ability to resolve pronoun ambiguity and encourages the development of models that understand the broad context in natural language text.
CoQA [316]
A conversational question-answering dataset, CoQA challenges models with questions that rely on conversation history and require free-form text answers. Its diverse content from seven domains makes it a rigorous test for models’ ability to handle a wide range of topics and conversational contexts.
WiC [317]
This dataset assesses a model’s ability to discern word meanings based on context, aiding in tasks related to Word Sense Disambiguation.
Wikitext103 [318]
With over 100 million tokens from Wikipedia’s top articles, this dataset is a rich resource for tasks that require understanding long-term dependencies, such as language modeling and translation.
PG19 [319]
This is a digital library of diverse books from Project Gutenberg. It’s specifically designed to facilitate research in unsupervised learning and language modeling, with a special focus on long-form content.
C4 [10]
A clean, multilingual dataset, C4 offers billions of tokens from web-crawled data. It’s a comprehensive resource for training advanced Transformer models on various languages.
LCQMC [320]
The Large-scale Chinese Question Matching Corpus (LCQMC) is a dataset for evaluating the performance of models in semantic matching tasks. It contains pairs of questions in Chinese and their matching status, making it a valuable resource for research in Chinese language understanding.
VI-B3 Story Cloze and Sentence Completion
StoryCloze [334]
It introduces a new “StoryCloze Test”, a commonsense reasoning framework for evaluating story understanding, generation, and script learning. It considers a model’s ability to understand and generate coherent and sensible stories.
LAMBADA [335]
This dataset evaluates contextual text understanding through a word prediction task. Models must predict the last word of a passage, which is easy for humans when given the whole passage, but not when given only the last sentence.
VI-B4 Physical Knowledge and World Understanding
PIQA [340]
A dataset that probes the physical knowledge of models, aiming to understand how well they are learning about the real world.
TriviaQA [341]
A dataset that tests models on reading comprehension and open domain question answering (QA) tasks, with a focus on Information Retrieval (IR)-style QA.
ARC [342]
A larger version of the ARC-Challenge, this dataset contains both easy and challenging grade-school level, multiple-choice science questions. It’s a comprehensive test of a model’s ability to understand and answer complex questions.
ARC-Easy [342]
A subset of the ARC dataset, ARC-Easy, contains questions that are answered correctly by either a retrieval-based algorithm or a word co-occurrence algorithm. It’s a great starting point for models beginning to explore advanced question-answering.
ARC-Challenge [342]
A rigorous question-answering dataset, ARC-Challenge includes complex, grade-school level questions that demand reasoning beyond simple retrieval, testing the true comprehension capabilities of models.
VI-B5 Contextual Language Understanding
RACE [347]
The RACE is a reading comprehension dataset collected from English examinations in China, which benchmarks AI models for understanding and answering questions on long and complex passages, simulating the challenge of a real-world examination.
RACE-Middle [347]
Another subset of the RACE [347] dataset, RACE-Middle, contains middle school-level English exam questions. It offers a slightly less challenging but academically oriented evaluation of a model’s comprehension skills.
RACE-High [347]
A subset of the RACE [347] dataset, RACE-High consists of high school-level English exam questions. It is designed to evaluate the comprehension ability of models in a more academic and challenging context.
QuAC [348]
This dataset simulates an information-seeking dialog between students and teachers using hidden Wikipedia text. It introduces unique challenges not found in machine comprehension datasets, making it a valuable resource for advancing dialog systems.
VI-B6 Commonsense Reasoning
HellaSwag [355]
A dataset that challenges models to pick the best ending to a context uses Adversarial Filtering to create a ‘Goldilocks’ zone of complexity, where generated text is absurd to humans but often misclassified by models.
COPA [402]
This dataset evaluates a model’s progress in open-domain commonsense causal reasoning. Each question comprises a premise and two alternatives, and the model must select the more plausible alternative, testing a model’s ability to understand and reason about cause and effect.
WSC [357]
The Winograd Schema Challenge (WSC) is a reading comprehension task in which a system must resolve references in a text, often requiring world knowledge and reasoning about the text.
CSQA [358]
The CommonsenseQA is a question-answering dataset that requires commonsense knowledge to answer the ability of AI models to understand and answer questions that require commonsense reasoning.
VI-B7 Reading Comprehension
BoolQ [363]
A dataset derived from Google search queries, BoolQ challenges models to answer binary (yes/no) questions. The questions are naturally occurring and are paired with a paragraph from a Wikipedia article containing the answer. It’s a test of reading comprehension and reasoning.
SQUADv2 [364]
The Stanford Question Answering Dataset (SQuAD) [362] is a collection of questions posed by crowdworkers on a set of Wikipedia articles, where the answer to every question is a segment of text from the corresponding reading passage. SQuADv2 combines the original SQuAD1.1 dataset with over 50,000 unanswerable questions. The aim is to evaluate a model’s ability to understand and answer questions based on a given context and to determine when a question is unanswerable.
DROP [365]
DROP, or Discrete Reasoning Over the content of Paragraphs, is designed to test a model’s ability to understand a wide variety of reading phenomena. It encourages comprehensive and reliable evaluation of reading comprehension capabilities.
RTE [366]
The Recognizing Textual Entailment (RTE) datasets come from a series of annual competitions on textual entailment, predicting whether a given sentence logically follows from another and evaluating a model’s understanding of logical relationships in a text.
WebQA [367]
A dataset for open-domain question answering, WebQA offers a large collection of web-based question-answer pairs. It is designed to assess the ability of AI models to understand and answer questions based on web content.
CMRC2018 [369]
This dataset is a test of Chinese language models’ ability to reason comprehensively and is designed with a challenging span-extraction format that pushes the boundaries of machine performance.
VI-B8 Mathematical Reasoning
MATH [382]
This dataset is a platform for evaluating the mathematical problem-solving abilities of AI models. It contains a diverse set of math problems, ranging from arithmetic to calculus, and is designed to test the model’s ability to understand and solve complex mathematical problems.
Math23k [383]
This one challenges a model’s ability to understand and solve mathematical word problems. It contains 23,000 Chinese arithmetic word problems that require models to perform reasoning and computation based on the problem description.
GSM8K [384]
A dataset of diverse grade school math word problems, testing a model’s ability to perform multi-step mathematical reasoning.
VI-B9 Problem Solving and Logical Reasoning
ANLI [394]
A large-scale dataset designed to test the robustness of machine learning models in Natural Language Inference (NLI) is created through an iterative, adversarial process where humans try to generate examples that models cannot correctly classify.
HumanEval [391]
A dataset for the problem-solving ability of AI models, which includes a diverse set of tasks that require various cognitive abilities, makes it a comprehensive tool for assessing general intelligence in AI.
StrategyQA [349]
A question-answering dataset that requires reasoning over multiple pieces of evidence to evaluate the strategic reasoning ability of AI models, pushing the boundaries of what machines can understand and answer.
VI-B10 Cross-Lingual Understanding
XNLI [399]
A cross-lingual benchmark, XNLI extends the MultiNLI [431] corpus to 15 languages, including low-resource ones like Urdu. It tests models on cross-lingual sentence understanding, with 112,500 annotated pairs across three categories: entailment, contradiction, and neutral.
PAWS-X [400]
PAWS-X, or Cross-lingual Paraphrase Adversaries from Word Scrambling, is a multilingual version of the PAWS [432] dataset for paraphrase identification. It includes examples in seven languages and is designed to evaluate the performance of cross-lingual paraphrase identification models.
VI-B11 Truthfulness
Truthful-QA [406]
A unique benchmark that measures a language model’s truthfulness when generating answers. The dataset includes questions across various categories like health, law, and politics, some designed to test the model against common human misconceptions.
VI-B12 Biases and Ethics in AI
ETHOS [409]
ETHOS is a hate speech detection dataset built from YouTube and Reddit comments. It’s a tool in the fight against online hate speech, offering binary and multi-label variants for robust content moderation.
StereoSet [410]
StereoSet is a comprehensive dataset designed to measure and evaluate the presence of stereotypical biases in language models. It focuses on four key domains: gender, profession, race, and religion. Contrasting stereotypical bias against language modeling ability provides a valuable tool for understanding and mitigating biases in large language models.
TABLE XII: Performance comparison of top performing LLMs across various NLU and NLG tasks. Here, “N-Shots” indicate the number of example prompts provided to the model during the evaluation, representing its capability in few-shot or zero-shot learning settings, “f” represents the fine-tuned version, and “B” represents the benchmark.
Task Dataset/Benchmark Model Model Size N-Shots Score
Multi-Task BIG-bench (B) Chinchilla 70B 5-shot 65.1
Gopher 280B 5-shot 53.97
PaLM 540B 5-shot 53.7
MMLU (B) GPT-4 - 5-shot 86.4
Gemini Ultra 5-shot 83.7
Flan-PaLM-2${}_{(f)}$ Large 5-shot 81.2
Language Understanding SuperGLUE (B) ERNIE 3.0 12B - 90.6
PaLM${}_{(f)}$ 540B - 90.4
T5 11B - 88.9
Story Comprehension and Generation HellaSwag GPT-4 - 10-shot 95.3
Gemini Ultra 10-shot 87.8
PaLM-2 Large one shot 86.8
StoryCloze GPT3 175B few shot 87.7
PaLM-2 Large one shot 87.4
OPT 175B - 79.82
Physical Knowledge and World Understanding PIQA PaLM-2 Large one shot 85.0
LLaMa 65B zero shot 82.8
MT-NLG 530B zero shot 81.99
TriviaQA PaLM-2 Large one shot 86.1
LLaMA-2 70B one shot 85.0
PaLM 540B one shot 81.4
Contextual Language Understanding LAMBADA PaLM 540B few shot 89.7
MT-NLG 530B few shot 87.15
PaLM-2 Large one shot 86.9
Commonsense Reasoning WinoGrande GPT-4 - 5-shot 87.5
PaLM-2 Large one shot 83.0
PaLM 540B zero shot 81.1
SIQA LLaMA 65B zero shot 52.3
Chinchilla 70B zero shot 51.3
Gopher 280B zero shot 50.6
Reading Comprehension BoolQ PaLM${}_{(f)}$ 540B - 92.2
T5 11B - 91.2
PaLM-2 Large one shot 90.9
Truthfulness Truthful-QA LLaMA 65B - 57
Mathematical Reasoning MATH Gemini Ultra 4-shot 53.2
PaLM-2 Large 4-shot 34.3
LLaMa-2 65B 4-shot 13.5
GSM8K GPT-4 - 5-shot 92.0
PaLM-2 Large 8-shot 80.7
U-PaLM 540B - 58.5
Problem Solving and Logical Reasoning HumanEval Gemini${}_{(f)}$ Ultra zero shot 74.4
GPT-4 - zero shot 67.0
Code Llama 34B zero shot 48.8
VII Applications
The application of Large Language Models (LLMs) has become a hot topic in both AI-related research communities and industries, with many emerging uses being discovered and explored daily. These models, capable of understanding and generating human-like text, have found meaningful applications across a variety of fields. This section will provide an overview of LLMs’ applications in medicine, education, science, mathematics, law, finance, robotics, and coding. While each of these domains poses different challenges, LLMs have opportunities to make significant contributions given their wide-ranging scope of applicability.
General Purpose: One of the most significant applications of LLMs is as a general-purpose tool for a wide variety of tasks, even those that have not been specifically trained for [433]. This is due to their inherent ability to understand, generate, and manipulate human-like text in a contextually relevant manner. This allows them to perform tasks ranging from simple language translation and question-answering to more complex tasks like summarization, text generation, and even programming help [434]. The utility of LLMs is further enhanced by their ability to adapt to the specific style and tone of the text they are processing, making the outputs more user-friendly and context-aware. In everyday applications, LLMs can be used as personal assistants, helping users draft emails or schedule appointments [435]; they can also be deployed in customer service to handle common questions, freeing up human resources for more complex issues; or applied to generate content for digital platforms like websites, by creating human-like text based on given prompts [436]. Moreover, LLMs play a crucial role in data analysis, where they can filter large volumes of text data, summarize key points, and find patterns that would take humans much longer time to identify [437]. Despite their wide-ranging applications, it is essential to remember that LLMs, like any AI system, are only as good as the data they have been trained on. We should always use LLMs carefully since they can unintentionally reproduce biases in their training data, leading to potentially unfair or inaccurate results.
Medicine: The application of LLMs in the field of medicine is reshaping healthcare delivery and research. For example, LLMs are increasingly used in clinical decision support systems to provide physicians with evidence-based treatment recommendations [438, 439, 440]. By analyzing patient data and medical literature, they can help identify potential diagnoses, suggest appropriate tests, and recommend optimal treatment strategies. Moreover, LLMs can also enhance patient interactions with healthcare systems; e.g., they can be used in chatbot applications [441, 442, 443] to answer patient queries about symptoms or medications, schedule appointments, and even provide essential health advice. For medical research, LLMs are used to extract and filter information from a considerable amount of medical literature, identify relevant studies, summarize findings, and even predict future research trends [444, 445, 446]. For medical education, LLMs can help create training materials, generate exam questions, provide detailed explanations of complex medical topics, and offer personalized feedback to students [447, 448, 449, 450]. They can also simulate patient interactions, enabling students to practice and improve their clinical skills. At a broader level, LLMs can assist in public health initiatives by analyzing media data to detect disease outbreaks, monitor public sentiment towards health policies, and disseminate health information in a clear and understandable manner [451]. When employing LLMs to support public health initiatives, addressing related issues such as data privacy, the necessity for explainability, and the potential risk of propagating biases [452, 453].
Education: The integration of LLMs into the educational sector offers opportunities to enhance learning experiences, teacher support, and educational content development. For students, by analyzing their learning styles, performance, and preferences, LLMs can provide customized study materials and practice questions to develop personalized learning experiences [454]. For teachers, LLMs can help to create lesson plans and grade assignments and generate diverse and inclusive educational content, significantly saving more time for teaching and student interaction [455, 456]. In language learning, LLMs serve as advanced conversational partners capable of simulating conversations in multiple languages, correcting grammar, enhancing vocabulary, and aiding pronunciation for the needs of fluency in practice [457]. Furthermore, LLMs improve accessibility in education by providing support for students with disabilities. They can generate real-time transcriptions for the hearing impaired, offer reading assistance for the visually impaired, and simplify complex texts for those with learning disabilities [453]. As LLMs continue to evolve, their applications in education can benefit more students and teachers from different perspectives in practice.
Science: Similar to medical applications, LLMs can expedite the research process by quickly analyzing and summarizing scientific literature. By briefing comprehensible and accessible research summaries, LLMs can assist researchers in staying up-to-date with the latest findings, even in fields outside their area of expertise [458, 459]. In addition, LLMs can aid scientists in formulating new hypotheses and research questions since their ability to process large-scale datasets allows them to unveil insights that might not be immediately apparent to human researchers [460]. Moreover, for scientific writing, LLMs can help researchers draft documents, suggest improvements, and ensure adherence to specific formatting guidelines [461, 462]. This not only saves time but also improves the clarity of scientific communication, enabling interdisciplinary teams to work together more effectively.
Maths: In addition to providing mathematical research and education support, LLMs can assist in solving mathematical problems by giving step-by-step explanations and guiding users through complex proofs and calculations. They can help identify errors in reasoning or computation and suggest corrections, serving as an invaluable tool for both learning and verification purposes [463, 464]. LLMs can be employed to check the validity of mathematical proofs, offering a preliminary filter before human review. While they are not a substitute for the meticulous work of mathematicians, they can help simplify the process of proof verification [465, 466]. Moreover, LLMs enhance accessibility to mathematics by translating complex concepts and findings into understandable language for non-specialists [467], where the gap between theoretical mathematics and applied contexts such as physics, engineering, and economics can be bridged.
Law: LLMs can assist with the thematic analysis of legal documents, including generating initial coding for datasets, identifying themes, and classifying data according to these themes. This collaborative effort between legal experts and LLMs has proved to be effective in analyzing legal texts such as court opinions on theft, improving both the efficiency and quality of the research [468]. Additionally, LLMs have been evaluated for their ability to generate explanations of legal terms, focusing on improving factual accuracy and relevance by incorporating sentences from case law. By feeding relevant case law into the LLM, the augmented models can generate higher-quality explanations with less factually incorrect information [469]. Moreover, LLMs can be trained with specialized domain knowledge to perform legal reasoning tasks [470] and answer legal questions [471].
Finance: LLMs like BloombergGPT [141], trained on extensive proprietary financial datasets, exhibit superior performance on financial tasks. This indicates the value of domain-specific training in creating LLMs that can more accurately understand and process industry-specific language and concepts. The introduction of FinGPT [472] as an open-source model offers transparent and accessible resources to develop novel applications such as robo-advising, algorithmic trading, and low-code solutions, ultimately expanding the capabilities of financial services. Both BloombergGPT and FinGPT show the adaptability of LLMs to the financial domain, with the former showing the power of custom datasets and the latter emphasizing a data-centric approach and low-rank adaptation techniques for customization. Moreover, LLMs demonstrate an ability to break down complex financial tasks into actionable plans, enabling end-to-end solutions that were previously unfeasible with a single model [473].
Others: The application of LLM in coding is introduced in Section III-A2. For the utility of LLMs in robotics, please refer to Section III-D.
VIII Summary and Discussion
VIII-A Architecture
Due to the gigantic scale of LLMs, minor changes in architecture and training strategies have a big impact on performance and stability. Here, we summarize key architectural modules used in various LLMs, leading to better performance, reduced training time and memory, and better training stability.
Layer Normalization is found to have a significant effect on the performance and training stability of LLMs. Pre-norm, that is normalizing inputs rather than outputs, is more common among LLMs stabilizing the training [6, 126, 104]. BLOOM [13] and AlexaTM [122] utilize an additional layer normalization before embedding layer to stabilize the training of large-scale models, while the model’s zero-shot generalization ability can be negatively impacted [13]. However, another study [33] finds that pre-norm degrades fine-tuned model performance as compared to post-norm, and there are no stability benefits of pre-norm beyond the 100B scale. Therefore, GLM-130B [33] used deep-norm which is a variant of post-norm for better downstream task performance after fine-tuning.
Positional Encoding effect performance and training stability of LLMs like other building blocks of a model. BLOOM [13] finds ALiBi outperforming learned and rotary positional encodings. Contrary to this, GLM-130B [33] identifies rotary positional encoding better than ALiBi. So, there is no conclusion in literature about the positional encodings yet.
Parallel Attention where attention and feed-forward layers are parallel to each other rather than sequential in transformer block has shown to reduce training time by 15%. There is no evidence of performance drop due to this change in literature and used by the models PaLM [15], GPT-NeoX [115], and CodeGen [130].
Multi-Query Attention has shared key and value attention heads in a transformer block while query attention heads are projected as usual. This reduces memory usage and speeds up sampling in autoregressive decoding. No performance degradation has been observed with this change and makes the training efficient allowing larger batch sizes. Multi-query attention is used in [15, 132].
Mixture of Experts allows easily scaling model to trillion of parameters [129, 118]. Only a few experts are activated during the computation making them compute-efficient. The performance of MoE models is better than the dense models for the same amount of data and requires less computation during fine-tuning to achieve performance similar to the dense models as discussed in [118]. MoE architectures are less prone to catastrophic forgetting, therefore are more suited for continual learning [129]. Extracting smaller sub-models for downstream tasks is possible without losing any performance, making MoE architecture hardware-friendly [129].
Sparse vs Dense Activated GPT-3 [6] uses sparse transformers [63] whereas GLaM [118] and PanGu-$\sum$ [129] use MoE [119] architecture to lower computational costs and increase the model size and capacity. According to the literature, sparse modules do not degrade the model’s performance [63]. However, more experiments are required to verify this statement.
VIII-B Training Strategies
Training models at a huge scale require some tricks to reduce training costs, avoid loss divergence and achieve better performance. We summarize and discuss some of these key tricks used in different LLMs.
Mixed Precision is a famous method for LLMs to reduce memory usage and improve training efficiency. In mixed precision, forward and backward passes are performed in FP16 format whereas optimizer states and master weights are kept in FP32 format [474]. A drawback associated with this format change is training instability due to a smaller value range resulting in loss spikes [33]. An alternative to FP16 is BF16 which has a comparatively larger range and performs some precision-sensitive operations like gradient accumulation and softmax in FP32 [13]. BF16 has better performance and training stability but uses more memory and is supported on specific hardware, for example, A100 GPUs. Therefore, its adoption in LLMs is limited.
Training Instability is a common issue in LLMs where loss divergence or spiking is observed multiple times during training. This happens in the presence of gradient clipping [15]. To mitigate this problem, many approaches suggest restarting training from an earlier checkpoint [15, 33, 118], skipping 200-500 earlier data batches at the point of divergence in [15] and re-shuffling batches in [118]. The embedding layer gradient shrink proves to further stabilize the training as its gradient norm is significantly larger than the other layers [33]. Another suggestion to improve training stability for larger models is not to use biases in dense and norm layers as in [15].
Weight Initialization plays a significant role in model convergence and training stability. GPT-NeoX [115] initializes feed-forward layers before residuals with $\frac{2}{L\sqrt{d}}$ as in [144] and other layers with small initialization scheme [475]. This avoids activations growing exponentially with the increasing depth. MT-NLG [114] found higher variance for weight initialization leads to unstable training, hence validating small initialization scheme [475]. Various models perform random weight initialization which can cause bad initialization, Galactica [138] suggests a longer warmup to negate the effect.
Learning Rate is important for stable training. It is suggested to use a lower value [13, 15, 124] with warmup and decay (cosine or linear). Usually, the learning rate is within the range $1e^{-4}$ to $8e^{-4}$. Moreover, MT-NLG (530B) [114] and GPT-NeoX (20B) [115] suggest interpolating learning rates based on the model size using the GPT-3 [6] models ranging between 13B and 175B. This avoids tuning the learning rate hyperparameter.
Training Parallelism 3D parallelism, a combination of data, pipeline and tensor parallelism, is the most utilized training parallelism approach in LLMs [33, 15, 14, 13, 114, 112, 109]. In addition to the 3D parallelism, BLOOM [13] uses zero optimizer [37] to shard optimizer states. PanGu-$\alpha$ [104] and PanGu-$\Sigma$ [129] go beyond the 3D parallelism and apply 5D parallelism which additionally contains optimizer parallelism and rematerialization.
Mode Switching adds task-related tokens at the beginning of the text during training. These tokens refer to the natural language understanding and natural language generation tasks which are shown to improve the downstream task performance in [89, 124, 122]. During fine-tuning and inference, tokens are appended based on the downstream tasks.
Controllable Text Generation Generating credible and controlled text from a pre-trained model is challenging. GPT-3 [6] and other LLMs use in-context learning to control generated text. While in-context learning helps in controlling the generated text, ERNIE 3.0 Titan [35] suggests using adversarial loss to rank its generated text for credibility and soft prompts such as genre, topic, keywords, sentiment, and length for better control on generated text.
VIII-C Pre-Training vs Instruction Tuning
While pre-training is important for the generalization of LLMs, instruction-tuning improves the performance of LLMs further and makes them useable. Therefore, it is suggested to perform instruction fine-tuning of pre-trained LLMs to use them effectively [16, 18, 20, 93, 157].
VIII-D Supervised Models vs Generalized Models
Although generalized models are capable of performing diverse tasks with good performance they have not yet outperformed models trained in supervised settings. The supervised trained models are still state-of-the-art in various NLP tasks by a large margin as shown in [6, 15, 18].
VIII-E Zero-Shot vs Few-Shot
LLMs perform well in zero-shot and few-shot settings. But the performance difference between zero-shot and few-shot is large for pre-trained models [6, 15], naming LLMs as meta-learners [6]. LLMs zero-shot evaluations underperform unsupervised methods in neural machine translation [6]. The literature shows pre-training is not enough for good zero-shot performance [15, 16]. To improve the zero-shot performance the literature suggests using instruction fine-tuning that improves the zero-shot performance significantly and outperforms baselines. Instruction fine-tuning has also been shown to improve zero-shot generalization to unseen tasks. Another model Flan-PaLM [16] unlocks zero-shot reasoning with CoT training.
VIII-F Encoder vs Decoder vs Encoder-Decoder
Traditionally, these architectures perform well for different tasks, for example, encoder-only for NLU tasks, decoder-only for NLG, and encoder-decoder for sequence2sequence modeling. Encoder-only models are famous for smaller models such as Bert [7], RoBERTa [424], etc, whereas LLMs are either decoder-only [6, 115, 13] or encoder-decoder [10, 11, 122]. While decoder-only models are good at NLG tasks, various LLMs, PaLM [15], OPT [14], GPT-3 [6], BLOOM [13], LLaMA [147], are decoder-only models with significant performance gains on both NLU and NLG tasks. In contradiction to this, T5 [10] and UL2 [89] identify encoder-decoder models out-performing decoder-only models. In another study, PaLM [15] finds increasing the size of decoder-only models can reduce the performance gap between decoder-only and encoder-decoder architectures.
Although decoder-only architectures have become a trend for LLMs, many recently proposed approaches [89, 122] use mode-switching tokens in text with encoder-decoder architectures to enable task-specific modes. Similarly, CodeT5+ [34] uses an encoder-decoder architecture with multiple training objectives for different tasks, activating the encoder, decoder, or both according to the tasks. These variations in architecture and training objectives allow a model to perform well in different settings. Because of this dynamic configuration, the future of LLMs can be attributed to encoder-decoder architectures.
IX Challenges and Future Directions
LLMs such as GPT-4 and its predecessors have significantly advanced natural language processing. Nevertheless, they also bring along a set of challenges. The computational cost, adversarial robustness, and interpretability are among the technical challenges that are intrinsic to these models. Furthermore, as these models are scaled up to handle more complex tasks or to operate in more complex or dynamic environments, new challenges in scalability, privacy, and real-time processing emerge. On the frontier of foundational research, integrating multi-modality and the effectiveness of transfer learning are being keenly explored. Additionally, the continuous learning aspect of these models, which aims to have models that can adapt to new information over time, presents a fresh set of challenges. These challenges not only underscore the technical intricacies involved but also highlight the broader impact and the future trajectory of LLMs in real-world applications. The following sections delve into these challenges, shedding light on the ongoing and potential efforts to address them.
Computational Cost: Training LLMs requires extensive computational resources, which increases production costs and raises environmental concerns due to substantial energy consumption during large-scale training. Improved performance occurs as computational resources increase, but the rate of improvement gradually decreases when both the model and dataset size remain fixed, following the power law of diminishing returns [476].
Bias and Fairness: LLMs can inherit and amplify societal biases in their training data. These biases can manifest in the model’s outputs, leading to potential ethical and fairness issues [477].
Overfitting: Although LLMs possess substantial learning capabilities, they are susceptible to overfitting noisy and peculiar patterns within their extensive training data. Consequently, this may cause them to generate illogical responses [478]. The debate about Memorization vs. Generalization in LLMs is about finding the right balance. Memorization allows the model to remember specific details from its training data, ensuring it can provide accurate answers to precise questions. However, generalization enables the model to make inferences and produce responses for inputs it hasn’t seen before, which is essential for handling various real-world tasks. Striking the right balance is the challenge: too much memorization can lead to overfitting, making the model inflexible and struggling with new inputs [479].
Economic and Research Inequality: The high cost of training and deploying LLMs may make their development concentrated within well-funded organizations, potentially worsening economic and research inequalities in AI [480].
Reasoning and Planning: Some reasoning and planning tasks, even as seemingly simple as common-sense planning, which humans find easy, remain well beyond the current capabilities of LLMs evaluated using an assessment framework. This isn’t entirely unexpected, considering that LLMs primarily generate text completions based on likelihood and offer no solid guarantees in terms of reasoning abilities [481].
Hallucinations: LLMs exhibit "hallucinations," where they generate responses that, while sounding plausible, are incorrect or don’t align with the provided information [482]. The hallucination can be categorized into three categories.
  • Input-conflicting hallucination, wherein LLMs produce content that diverges from the input given by users.
  • Context-conflicting hallucination, where LLMs generate content that contradicts information they have generated earlier.
  • Fact-conflicting hallucination involves LLM’s generation of content that does not align with established world knowledge.
Prompt Engineering: Prompts serve as inputs to LLMs, and their syntax and semantics play a crucial role in determining the model’s output. The prompt variations, sometimes counter-intuitive to humans, can result in significant changes in model output and are addressed through prompt engineering, which involves designing natural language queries to guide LLMs responses effectively [483, 32].
Limited Knowledge: Information acquired during pretraining is limited and may become obsolete after some time. Re-training the model using updated data is costly. To generate factually accurate responses people use retrieval augmentation pipeline [243]. However, pre-trained models are not trained with retrieval augmentation generation (RAG) [6, 21], hence, adapting the training pipeline is necessary [238, 25].
Safety and Controllability: Using LLMs comes with the risk of generating harmful, misleading, or inappropriate content, whether by accident or when given specific prompts. Ensuring these models are safely utilized is a significant concern [484].
Multi-Modality: Multi-modal learning, where LLMs are trained on diverse data like text, images, and videos, aims to create models with richer understanding but faces challenges in data alignment, fusion strategies, and higher computational demands.
Catastrophic Forgetting: LLMs are often pre-trained on large datasets and then fine-tuned on domain-specific data, reducing training resources but facing issues like domain adaptation and catastrophic forgetting, which hinders the retention of original knowledge when learning new tasks.
Adversarial Robustness: Large Language Models (LLMs) have shown great capabilities in various tasks but are vulnerable to adversarial attacks, where slight, deliberate input alterations can mislead them. Especially with models like BERT, adversarial fine-tuning can enhance robustness, although it sometimes compromises generalization [485]. As LLMs integrate more into complex systems, examining their security properties becomes crucial, given the emerging field of adversarial attacks on LLMs within trustworthy ML [486]. This vulnerability is notable in safety-critical domains, necessitating robust adversarial evaluation tools to ensure LLM reliability [487].
Interpretability and Explainability: The "black-box" nature of LLMs poses challenges in understanding their decision-making, which is crucial for broader acceptance and trust, especially in sensitive domains. Despite their advanced capabilities, the lack of insight into their operation limits their effectiveness and trustworthiness [488, 489]. Efforts are being made to make LLMs more explainable to promote user trust and to ensure responsible AI usage. Understanding the logic behind LLMs’ responses is essential for fostering trust and ensuring they align with human values and legal standards.
Privacy Concerns: Privacy concerns in Large Language Models (LLMs) have escalated with their growth in complexity and size, particularly around data sharing and potential misuse. There is a risk of malicious content creation, filter bypass, and data privacy issues, especially in e-commerce, where protecting customer privacy is crucial. If models are trained on private data, additional concerns arise if such models are made publicly available. LLMs tend to memorize phrases from their training sets, which an adversary could exploit to extract sensitive data, posing a threat to personal privacy [490, 491].
Real-Time Processing: Real-time processing in Large Language Models (LLMs) is pivotal for various applications, especially with the rising popularity of mobile AI applications and concerns regarding information security and privacy. However, LLMs often have hundreds of layers and millions of parameters, which impede real-time processing due to the high computational demands and limited weight storage on hardware platforms, particularly in edge computing environments [492]. While certain efforts like MobileBERT aim to reduce memory requirements, they still face substantial execution overhead due to the large number of model layers, leading to high inference latency.
Long-Term Dependencies: Large Language Models (LLMs) have shown considerable progress in understanding and generating text, yet they often struggle with preserving context and handling long-term dependencies, particularly in complex, multi-turn conversations or long documents. This limitation can lead to incoherent or irrelevant responses.
Hardware Acceleration: The growth of LLMs presents significant hardware challenges due to the increasing computational and memory demands associated with training and deploying these models. GPUs have played a crucial role in meeting the hardware requirements for training LLMs, with the networking industry also evolving to optimize hardware for training workloads. However, the growing size of LLMs, which has been outpacing hardware progress, makes model inference increasingly costly. Model quantization is a promising approach to bridge the widening gap between LLM size and hardware capacity [493]. Although specialized hardware acceleration like GPUs or TPUs can significantly reduce the computational cost, making real-time applications more feasible, they may not fully resolve all limitations, necessitating further advancements in hardware technology.
Regulatory and Ethical Frameworks: The rapid advancements in artificial intelligence have given rise to sophisticated Large Language Models (LLMs) like OpenAI’s GPT-4 [148] and Google’s Bard. These developments underscore the imperative for regulatory oversight to manage the ethical and social challenges accompanying LLMs’ widespread use [494]. For instance, LLMs can generate content that can be used positively or negatively, emphasizing the need for proactive ethical frameworks and policy measures to guide their responsible use and assign accountability for their outputs [495]. Auditing is identified as a promising governance mechanism to ensure that AI systems, including LLMs, are designed and deployed ethically, legally, and technically robust [496].
X Conclusion
This article has reviewed various LLMs, discussing their interesting aspects. It contributes to summarizing significant findings in the existing literature and provides a detailed analysis of the design aspects of LLMs, including architectures, datasets, and training pipelines. We identified crucial architectural components and training strategies employed by different LLMs. These aspects are presented as summaries and discussions throughout the article. Moreover, we have discussed the performance differences of LLMs in zero-shot and few-shot settings, explored the impact of fine-tuning, and compared supervised and generalized models and encoder vs decoder vs encoder-decoder architectures. A comprehensive review of LLMs in robotics, multi-modal LLMs, augmented LLMs, datasets, and evaluation is also provided. This article is anticipated to serve as a valuable resource for researchers, offering insights into the recent advancements in LLMs and providing fundamental concepts and details to develop better LLMs.
XI Versioning
We keep track of the versions of this paper we release as the content updates.
Version 1.0: We covered 30 pre-trained models and 6 instruction-tuned models, including their overview, findings, training, and evaluation datasets, and discussed important architectural and training tricks by various LLMs.
Version 2.0: Further pre-trained LLMs added along with discussion on on self-instruct LLMs. Categorized LLMs according to the application, provided descriptions of widely used evaluation datasets, added a section on robotics, and extended discussion in section VIII. Tables have been updated.
Version 3.0: Added sections on Alignment tuning and multimodal LLMs. A performance comparison table on various benchmarks and datasets. Added LLaMA-2 and PaLM-2.
Version 4.0: Tables on training and evaluation datasets, a sub-section on increasing context window, and minor improvements.
Version 5.0: Added sections on augmented LLMs and challenges and future directions.
Version 6.0: Minor improvements in abstract, introduction, and conclusion.
Version 7.0: Sections on efficient LLMs and applications.
Note: If you find any mistakes, or have issues and conflicts with the writing in this paper, please email us. We welcome suggestions to improve this paper.